### v1

# General Time Series Forecasting

Abstract
In this work, we introduce FLAME, a family of extremely lightweight and capable Time Series Foundation Models, which support both deterministic and probabilistic forecasting via generative probabilistic modeling, thus ensuring both efficiency and robustness. FLAME utilizes the Legendre Memory for strong generalization capabilities. Through adapting variants of Legendre Memory, i.e., translated Legendre (LegT) and scaled Legendre (LegS), in the Encoding and Decoding phases, FLAME can effectively capture the inherent inductive bias within data and make efficient long-range inferences. To enhance the accuracy of probabilistic forecasting while keeping efficient, FLAME adopts a Normalization Flow based forecasting head, which can model the arbitrarily intricate distributions over the forecasting horizon in a generative manner. Comprehensive experiments on well-recognized benchmarks, including TSFM-Bench and ProbTS, demonstrate the consistent state-of-the-art zero-shot performance of FLAME on both deterministic and probabilistic forecasting tasks.
Time Series Forecasting
1 Introduction
Time Series Forecasting is a critical and widely studied task due to the attractive charm of peeking future, which helps make significant values in various fields through planning ahead. While, time series dances like a flame – it is hard to accurately capture the determinstic rules. Therefore, to control risks and make reasonable decisions, quantifying uncertainties through Probabilistic Time Series Forecasting (PTSF) is a more reliable approach, which is continously explored for decades from statistical approaches (Godahewa et al., 2021a; Box & Pierce, 1970) with solid theories and interpretability to data-driven approaches (Nie et al., 2023; Qiu et al., 2025e; Wu et al., 2025b) with strong fitting capability. Among them, deep learning models (Nie et al., 2023; Qiu et al., 2025e; Wu et al., 2025b) show outstanding impressive performance on in-distribution tasks through effectively capturing the inherent dynamics within raw data. Inspired by the success of pretrained models in other domains such as CV (Khan et al., 2022; Liu et al., 2021) and NLP (Liu et al., 2024a; Achiam et al., 2023) with strong generalization capabilities, a surge of pretrained time series foudation models (Liu et al., 2024c; Das et al., 2024; Gao et al., 2024; Shi et al., 2024; Wang et al., 2025b, c) are developed for out-of-distribution tasks, demonstrating promising trends.
Figure 1: FLAME is pretrained on Time Series Corpus from Multiple Domains, possessing a lightweight backbone while supporting generative probabilistic forecasting.
Recent studies on time series foundation models (Woo et al., 2024; Shi et al., 2024; Wang et al., 2025b; Liu et al., 2025d) have achieved some progress in exploiting the generalization capabilities and exhibit excellent forecasting performance in zero-shot scenarios. Their generalization capabilities mainly come from two perspectives–see Figure 1: 1) some works make efforts in devising appropriate Transformer-based architectures, and expanding the parameters to Hundreds-Million- or Billion-Scale (Shi et al., 2024; Ansari et al., ; Woo et al., 2024; Das et al., 2024), or with the help of the massive parameters from LLMs (Zhou et al., 2023; Jin et al., 2024) and ViTs (Chen et al., 2024a). Though these works achieve excellent performance in zero-shot scenarios, they consume a lot in inference under limited resources, and some of them are even slower than training end-to-end supervised models from scratch, thus hindering the efficiency and practicality; 2) another branch of works explore the effective capture of inductive bias within data, with merticulously-designed lightweight structures (Ekambaram et al., 2024; Gao et al., 2024; Wang et al., 2025b) but strong generalization capabilites. However, although they demonstrate strong performance in determinstic forecasting tasks, they lack the support of probabilistic forecasting. Moreover, predicting accurate distributions over the forecasting horizon is difficult because assuming prior distributions like Isotropic Gaussian Distributions constrains the modeling capabilites of real data distributions. On the other hand, modeling arbitrary complex distributions relies on generative mechanisms (Li et al., 2017; Kollovieh et al., 2023), which may increase the model complexity and hinder the efficiency.
As shown in Figure 1, our proposed FLAME pionners the exploration of lightweight time series foundation models capable of both determinstic and probabilistic forecasting. To enhance the generalization capability while keeping lightweight, we utilize the Legendre Memory (Voelker et al., 2019; Gu et al., 2020) in both Encoding and Decoding phases to capture the inherent inductive bias within data, which contributes to a lightweight but strong framework. Originating from ordinary differential equations (ODEs) and characterized by orthogonal legendre polynomials, the Legendre Memory has been utilized in control theory to model continuous time-delay systems (Richard, 2003) for several decades, which has the potential to compress time series of varying lengths into vectors of the same size. This helps ensuring lightweight, and we employ Legendre Memory in both the encoder and decoder.
First, in the pretraining process of time series foundation models, the inherent periods and trends of data from distinct domains can lead to complex local characteristics but the patching technique cannot handle this due to the fixed receptive field. This inspires us to devise a Local-Perception module to unlock the power of Legendre Memory in integrating intricate and varying local characteristics within data as the supplement environmental semantic information to patched tokens. Second, we also adopt the Structured State-Space Duality (SSD (Dao & Gu, 2024)) layers in the decoding phase, which utilizes the Legendre Memory to initialize the state transition matrices (Gu et al., 2020), exceling at modeling long time series based on its selective states and adaptive memory utilization.
Besides the adaptions of Legendre Memory, we also devise a Flow-based Head to support generative probabilistic forecasting. Specifically, we focus on token-wise modeling, which constructs the distributions over dozens of time series points at once, and using the Flow-based Head driven by normalization flow (Rasul et al., 2021b; Durkan et al., 2019) can preserve both considerable efficiency and interpretability compared with prevailing diffusion models (Kollovieh et al., 2023; Rasul et al., 2021a). Additionally, we also devise to capture the token-wise causality in the Flow-based Head, which is important but neglected in current decoding methods (Liu et al., 2025d, 2024c; Shi et al., 2024), and this can help preserve temporal dependencies within tokens to boost the fine-grained modeling. Through these adaptions, the Flow-based Head can preserve both the accuracy and efficiency through constructing lightweight couple layers to simulate the reversible conditional transfer process from fixed prior Gaussian distributions to any complex real-data distribution. With the Legendre Memory and Flow-based Head facilliating the strong generative paradigm, FLAME can make extreme fast and accurate probabilistic forecasts. Our contributions are summarized as:
  • We present FLAME, a family of Flow enhanced Legendre Memory Foundation Models, including three parameter versions, i.e., 2M, 6M, 10M. Even the largest version of FLAME (10M) is far more lightweight than other time series foundation models which supports both determinstic and probabilistic forecasting.
  • We adapt the Legendre Memory to enhance the generalization capability. Both the Local-Perception Module and SSD-Decoder are driven by Legendre Memory and able to adaptively utilize the inductive bias within data.
  • We devise the Flow-based Head to model the intricate token-wise distributions over the forecasting horizon, considering the causality within tokens, with strong efficiency and accuracy.
  • Experimentally, FLAME achieves state-of-the-art zero-shot performance on both determinstic and probabilistic forecasting benchmarks, including TSFM-Bench (Li et al., 2025c) and ProbTS (Zhang et al., 2024), demonstrating a strong out-of-the-box tool of decision intelligence. The resources are available at https://anonymous.4open.science/r/FLAME-D83F.
2 Legendre Memory
When modeling sequence data with recurrent networks, “memory” is a common concept denoting the online compression capability of historical observations. Among them, Legendre Memory (Voelker et al., 2019; Gu et al., 2020) is a kind of continuous online function approximation, which utilizes orthogonal Legendre polynomials to compress the continous-time history $f$ into state vectors $m$ with fixed sizes. From the perspective of Real Analysis, given a time-varying measure $\mu^{(t)}$ and a high-order polynomial operator, recent history $f$ can be compressed into state vectors $m$ under certain approximation error (Gu et al., 2020). When the operator is the Legendre polynomial, there exists two common time-varying measures: translated Legendre (LegT) and scaled Legendre measure (LegS).
LegT assigns uniform weights to the most recent history $t-\theta,t]$, where $\theta$ denotes the length of the sliding windows, or the length of history that is being summarized:
$\text{LegT}:\mu^{(t)}(x)=\frac{1}{\theta}\mathbb{I}_{[t-\theta,t]}(x)$ (1)
Under the LegT measure, the process of compressing the historical information $f$ into state vectors $m$ can be represented as solving $d$ coupled linear time-invariant ODEs:
$m^{\prime}(t)=-Am(t)+Bf(t),$ (2)
$B_{n}=\frac{1}{\theta}(2n+1)(-1)^{n},$ (3)
$A_{nk}=\frac{1}{\theta}\begin{cases}(-1)^{n-k}(2n+1)&if\ n>k\\ -2n+1&if\ n\leq k\end{cases}\ ,$ (4)
where $A\in\mathbb{R}^{d\times d},B\in\mathbb{R}^{d\times 1}$ are optimal parameters derived through the use of Padé approximant (Padé, 1892; Voelker, 2019). In the forward process, the sequence of $t-\theta,t]$ can be compressed into the memory cell $m(t)\in\mathbb{R}^{d}$ in a recurrent mode using discretized Formula (2):
$m_{t}=\bar{A}m_{t-1}+\bar{B}f_{t},$ (5)
$\bar{A}=I-(\Delta t/\theta)A,\bar{B}=(\Delta t/\theta)B,$ (6)
where $\bar{A}$ and $\bar{B}$ can be obtained through zero-order hold (ZOH). And the “memory” lies in the capability of using $m(t)$ to recover all the past history $f$ of $t-\theta,t]$, thus can describe a $\theta$ time-delay system (Richard, 2003):
$f(t-\theta^{\prime})\approx\displaystyle\sum_{i=0}^{d-1}\mathcal{P}_{i}\left(\frac{\theta^{\prime}}{\theta}\right)m_{i}(t),\ 0\leq\theta^{\prime}\leq\theta,$ (7)
$\mathcal{P}_{i}(r)=(-1)^{i}\sum_{j=0}^{i}\begin{pmatrix}i\\ j\end{pmatrix}\begin{pmatrix}i+j\\ j\end{pmatrix}(-r)^{j},$ (8)
where $\mathcal{P}_{i}(r)$ is the $i$-th shifted Legendre polynomial (Voelker et al., 2019). According to Theorem 2.1 (Gu et al., 2020), larger number $d$ of Legendre Polynomials basis leads to more accurate approximation.
Theorem 2.1.
If $F(x)$ is $L$-Lipschitz, then $\|F_{[t-\theta,t]}(x)-f_{[t-\theta,t]}(x)\|_{\mu^{(t)}}\leq\mathcal{O}(\theta L/\sqrt{d})$. Moreover, if $F(x)$ has $k$-th order bounded derivatives, we have $\|F_{[t-\theta,t]}(x)-f_{[t-\theta,t]}(x)\|_{\mu^{(t)}}\leq\mathcal{O}(\theta^{k}d^{-k+1/2})$.
Compared with LegT, LegS assigns uniform weights to all history $[0,t]$ to support the compression of the entire history:
$\text{LegS}:\mu^{(t)}(x)=\frac{1}{t}\mathbb{I}_{[0,t]}(x)$ (9)
Under the LegS measure, the compressing process is:
$m^{\prime}(t)=-\frac{1}{t}Am(t)+\frac{1}{t}Bf(t),$ (10)
$B_{n}=(2n+1)^{\frac{1}{2}},$ (11)
$A_{nk}=\begin{cases}(2n+1)^{1/2}(2k+1)^{1/2}&if\ n>k\\ n+1&if\ n=k\\ 0&if\ n<k\end{cases}\ ,$ (12)
where $A\in\mathbb{R}^{d\times d}$ and $B\in\mathbb{R}^{d\times 1}$ can also be discretized as Formula (5) and Formula (6).
Figure 2: The overall framework of FLAME. The data flow follows the sequence of (a) - (f).
3 FLAME
FLAME adopts the Channel-Independent (Nie et al., 2023) pretraining paradigm, and each variable is preprocessed through Instance Normalization (Ulyanov et al., 2016) to mitigate the value discrepancy. As shown in Figure 2, FLAME utilizes the Re-Norm (Kim et al., 2021) to further mitigate the statistical differences between inputs and forecasts, and its backbone mainly consists of three modules: 1) Encoding, including Time Series Tokenization, Local-Perception, and MSA-Encoder, which tokenize the time series and enhance them through fusing the local environmental information with LegT; 2) Decoding, including LegS based SSD-Decoder and MCA-Enhancer, which utilize the SSD layers and MCA layers to make long-term inference; 3) Flow-based Head, which leverages the Normalization Flow to support generative probabilistic forecasting.
3.1 Encoding
3.1.1 Time Series Tokenization
In the pretraining Time Series Corpus, all data are preprocessed into univariate time series and received by FLAME in a Channel-Independent paradigm. Following the commonly used patching technique, which is first introduced by Triformer (Cirstea et al., 2022), we first divide each univariate time series $X\in\mathbb{R}^{T}$ into non-overlapping time series patches:
$X^{P}=\left[h_{1},h_{2},\cdots,h_{n}\right],$ (13)
where $X^{P}\in\mathbb{R}^{n\times p}$ are the patches, $n=\lceil T/p\rceil$ denotes the patch numer and $p$ is the patch size. $h_{i}\in\mathbb{R}^{p}$ is the $i$-th patch. Subsequently, we use a linear layer: $\mathbb{R}^{p}\to\mathbb{R}^{d}$ to transform the patches to tokens:
$X^{token}=\left[\text{linear}(h_{1}),\text{linear}(h_{2}),\cdots,\text{linear}(h_{n})\right],$ (14)
where $X^{token}\in\mathbb{R}^{n\times d}$ are the embeded time series tokens. Patch-wise tokens preserve semantic information and reduce the number compared with point-based tokens, thus enhancing efficiency.
3.1.2 Local-Perception
Though patching technique (Cirstea et al., 2022; Nie et al., 2023) has achieved some progress in time series foundation models, its fixed receptive field may limit models’ generalization capabilities, thus requiring heavy backbones. The core reason is that the inherent periods and trends within data vary across datasets from distinct domains, causing the diversity of frequencies, but the fixed patch size in pretrained models only captures the same span without considering the changing surrounding environmental information.
Specifically, in pretraining scenarios, a patch with the fixed length may cover the whole period in one pretraining dataset and only covers, e.g., the 70% period in another–see Figure 2(b), which causes burden to understand this semantic transformation. To tackle this limitation, we manage to recognize the varying periods, then align their information with patch-wise tokens to supplement the environmental semantics. First, we adopt the FFT (Brigham & Morrow, 1967) to find the period of the input series:
$\mathcal{A}=\text{Amp}(\text{FFT}(X)),\mathcal{F}=\text{arg}\max(\mathcal{A}),\mathrm{P}=\left\lceil\frac{T}{\mathcal{F}}\right\rceil,$ (15)
where $\text{FFT}(\cdot)$ and $\text{Amp}(\cdot)$ denote the FFT and the calculation of amplitude values. $\mathrm{P}$ is the searched period length. Then we consider the relative positional relations between the patch and the surronding period to make alignment:
$s=\lfloor(p+\mathrm{P}-1)/\mathrm{P}\rfloor\cdot\mathrm{P}-p,$ (16)
$X^{\prime}=\text{LeftPadding}(X,s),$ (17)
$X^{P^{\prime}}=[w_{1},w_{2},\cdots,w_{n}],$ (18)
where $X^{P^{\prime}}\in\mathbb{R}^{n\times(p+s)}$ are the local environmental patches. The padding length $s$ is used to make each local environmental patch $w_{i}\in\mathbb{R}^{p+s}$ cover the corresponding patch $h_{i}\in\mathbb{R}^{p}$, and $p+s$ is the integer multiple of the period $\mathrm{P}$. These operations can help align the intricate and varying local information $w_{i}$ to the patch $h_{i}$.
For each $w_{i}\in\mathbb{R}^{p+s}$, its length varies with the sample during pretraining due to the multiple frequencies of the time series corpus, which is hard to be encoded through predefined neural networks like MLPs. To overcome this, we propose to utilize the LegT Legendre Memory–see Section 2, to encode the varying local environmental information, i.e., the varying periods, into vectors of the fixed size. For each $w_{i}\in\mathbb{R}^{p+s}$, we use the discretized form of LegT to compress it like a state space model in an autoregressive manner:
$\bar{A}=I-(\Delta t/\theta)A,\bar{B}=(\Delta t/\theta)B,$ (19)
$f_{t}=w_{i,t},m_{0}=0,m_{t}=\bar{A}m_{t-1}+\bar{B}f_{t},$ (20)
where $\theta=p+s$ is determined based on the length of $w_{i}$, $A\in\mathbb{R}^{d\times d}$ and $B\in\mathbb{R}^{d\times 1}$ are untrainable and initialized through Formula (4), $\Delta t$ is a predefined parameter to control the granularity of zero-order hold (ZOH). This mechanism compresses the local environmental patch $w_{i}\in\mathbb{R}^{p+s}$ with varying lengths into the memory cell $m_{t}\in\mathbb{R}^{d}$ with the fixed size. The generated environmental tokens are:
$X^{token^{\prime}}=[w^{\prime}_{1},w^{\prime}_{2},\cdots,w^{\prime}_{n}],$ (21)
where each token $w^{\prime}_{i}\in\mathbb{R}^{d}$ is the last state of memory cell. We then fuse the original tokens $X^{token}$ and the environmental tokens $X^{token^{\prime}}$ into $H\in\mathbb{R}^{n\times d}$: $H=X^{token}+X^{token^{\prime}}$. According to Theorem 2.1, the dimension $d$ is often greater than $\theta$, which helps preserve the full semantics.
3.1.3 MSA-Encoder
After obtaining the fused tokens, we use the Multi-head Self-Attention Layer as the encoder to generate the hidden representations of them. Through fully-connected interaction, the MSA-Encoder aims at learning the similarities and differences among patterns and environmental information within the tokens:
$H^{\prime}=\text{MSA-Encoder}(H),$ (22)
the representations $H^{\prime}\in\mathbb{R}^{n\times d}$ are the outputs of the Encoding phase, and provide rich semantics within data.
3.2 Decoding
3.2.1 SSD-Decoder
We utilize LegT in Local-Perception of the Encoding phase to enrich the local environmental information. In the Decoding phase, the long-range modeling capability gains more attention, which calls for adaptive memory utilization when making long-term forecasts. Structured State-Space Duality (SSD) Layer or Mamba2 (Dao & Gu, 2024), uses the selective state space modeling mechanism to adaptively process historical information, significantly improving efficiency while achieving similar accuracy to Transformers. Its mechanism takes the form as follows:
$m_{t}=A_{t}m_{t}+B_{t}x_{t},y_{t}=C^{T}_{t}m_{t},$ (23)
where the form shares similarities with LegS, see–Formula (10). Previous works (Gu & Dao, 2023; Dao & Gu, 2024) demonstrate that adapting $A,B,C$ to selective $A_{t},B_{t},C_{t}$ can enhance the non-linearity capability. Among them, $A_{t}$ is obtained from $A$ through the learnable granularity $\Delta t$ of discretization, so that the initialization of $A$ still matters. To preserve the capability of adaptive memory utilization and ensure the linear complexity of SSD, we use the diagnoal of $A$ in LegS–see Formula (12), as the initialization of $A_{t}$. We then use the SSD layer with initialized $A_{t}$ as the Decoder to forecast the future tokens:
$F=\text{SSD-Decoder}(H^{\prime}),$ (24)
where $F\in\mathbb{R}^{l\times d}$. Supposing the length of forecasting horizon is $L$, $l=\lceil L/p\rceil$. Based on the LegS, SSD-Decoder can efficiently memorize all the information from inputs and adaptively utilize them like Transformers.
3.2.2 MCA-Enhancer
We then use the Multi-head Cross-Attention Layer to further enhance the outputs through recalling the similar representations in the inputs. Specifically, we set $F$ as Query, $H^{\prime}$ as Key and Value. Through skip connections in MCA-Enhancer, it effectively enhances the $F$:
$F^{\prime}=\text{MCA-Enhancer}(F),$ (25)
where $F^{\prime}\in\mathbb{R}^{l\times d}$ are the outputs of the Decoding phase, which contain the high-order semantic information obtained from the token transition process in the hidden spaces.
3.3 Flow-based Head
Through pretraining on Time Series Corpus from multiple domains, the Encoding and Decoding phases of FLAME possess strong capability of modeling the token transition principles, which means $F^{\prime}\in\mathbb{R}^{l\times d}$ are high-quality representations for forecasting. In time series foundation models, making token-wise forecasts through relative lightweight heads can preserve efficiency and ensure the heavy backbones do learn the “knowledge” (Liu et al., 2024c, 2025d).
However, since the inherent temporal causality within time series data, patch-based forecasting models only model the token-wise causal relationships in decoding. When transforming these high-dimension tokens into patch sizes for forecasts, the fine-grained causal relationships within each patch are often neglected. To support generative probabilistic forecasting, keep lightweight, and capture such fine-grained causal relationships, we merticulously devise the couple layers of Normalization Flow to model the intricate probabilistic distributions of real-data. The whole generative process is fourmulated as:
$F^{\prime}=[o_{1},o_{2},\cdots,o_{l}],$ (26)
$\epsilon_{i}\sim\mathcal{N}(0,I),\mathcal{F}_{i}=\text{Concat}[\epsilon_{i},o_{i}],$ (27)
$\hat{Y}_{i}=\ f_{K}\circ f_{K-1}\circ\cdots\circ f_{1}(\mathcal{F}_{i}),$ (28)
where we utilize the tokens $F^{\prime}=[o_{1},o_{2},\cdots,o_{l}]\in\mathbb{R}^{l\times d}$ as the prior conditions, and the Gaussian Noise $\epsilon_{i}\in\mathbb{R}^{p}$ is the initialization. We concat them to obtain $\mathcal{F}_{i}\in\mathbb{R}^{p+d}$ as the input of couple layers. The structure of couple layers $f_{i}$ is the masked MLP:
$\mathcal{M}_{n,k}=\begin{cases}0&if\ \ n<k\leq p\\ 1&otherwise\\ \end{cases},$ (29)
$\mathcal{F}^{\prime}_{i}=\ f_{i}(\mathcal{F}_{i})=\text{MLP}^{i}_{\mathcal{M}}(\mathcal{F}_{i}),$ (30)
where the mask $\mathcal{M}\in\mathbb{R}^{(p+d)\times(p+d)}$ determines the connection pattern of neurons in MLP-based couple layer $f_{i}:=\text{MLP}^{i}_{\mathcal{M}}$, which determines the interactions among dimensions of $\mathcal{F}_{i}\in\mathbb{R}^{p+d}$. Considering the output $\mathcal{F}^{\prime}_{i}\in\mathbb{R}^{p+d}$, $\mathcal{M}_{n,k}=0$ indicates that $\mathcal{F}^{\prime}_{i}[k]$ cannot see $\mathcal{F}_{i}[n]$ when calculated in $f_{i}$, and vice versa. This means $f_{i}$ allows the conditional part $o_{i}\in\mathbb{R}^{d}$ of $\mathcal{F}_{i}$ to self-interact and interact with all the dimension of $\epsilon_{i}\in\mathbb{R}^{d}$, and works in an autoregressive manner to ensure the noise part $\epsilon_{i}[j]$ only interacts with $\epsilon_{i}[\leq j]$ to preserve the fine-grained causal relationships within each patch–see Figure 2(f). Through $K$ coupled layers of generative probabilistic modeling, the final outputs are:
$\hat{Y}=[\hat{Y}_{1},\hat{Y}_{2},\cdots,\hat{Y}_{l}],$ (31)
where $\hat{Y}\in\mathbb{R}^{l\times p}$ can be further flattened and fetched the first $L$ points as the forecasts. With the MLP-based couple layers, Flow-based Head can achieve intricate distributional transformation while keeping lightweight. Due to the probabilistic generative mechanism, multiple samplings are recommended to evaluate the forecasted distributions. For clarity, we replace $\hat{Y}_{i}$ and $\mathcal{F}_{i}$ with $x$ and $z$ in the token-wise optimization objective:
$\log p_{X}(x)=\log p_{Z}(z)-\sum_{k=1}^{K}\log\left|\det\left(\frac{\partial f_{k}(z_{k-1})}{\partial z_{k-1}}\right)\right|$ (32)
4 Experiments
We introduce the detailed experimental settings in Appendix A, and show the zero-shot performance of FLAME in Section 4.1 (deterministic) and Section 4.2 (probabilistic). We also make detailed analysis on key components of FLAME to reveal where the generalization capability comes in Section 4.3. To analyze the scability of FLAME, we make studies on training and inference in Section 4.4. As an important property for deployment, we also analyze the efficiency of FLAME in Section 4.5. In summary–see Figure 3, our proposed FLAME achieves most 1st counts and keeps the most lightweight. We provide all results in Appendix C, including full results of Table 1 and 2, and other results on full-shot settings and short-term forecasting.
Figure 3: The parameter scale and zero-shot performance of Time Series Foundation models. FLAME Large achieves most 1st counts while keeping the most lightweight.
Table 1: Average results of zero-shot deterministic forecasting experiments on datasets from TSFM-Bench. Lower MSE or MAE values indicate better predictions. (’-’) denotes datasets included in the model’s pretraining and therefore excluded from testing. Red: the best, Blue: the 2nd best. All the results are listed in Table 13 of Appendix C.
Models FLAME FLAME FLAME Sundial LightGTS ROSE Timer Chronos Time-MoE TinyTTM MOIRAI
Small Base Large (2025) (2025) (2025) (2024) (2024) (2024) (2024) (2024)
Metrics MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETT (Avg) 0.339 0.382 0.339 0.381 0.344 0.385 0.335 0.379 0.339 0.379 0.393 0.411 0.551 0.478 0.442 0.408 0.357 0.390 0.441 0.430 0.382 0.388
Weather 0.215 0.261 0.222 0.269 0.227 0.273 0.234 0.270 0.219 0.267 0.265 0.305 0.292 0.313 0.288 0.309 0.256 0.289 0.265 0.307 0.260 0.275
Electricity 0.287 0.371 0.222 0.319 0.217 0.313 0.169 0.265 0.233 0.319 0.234 0.320 0.297 0.375 - - - - 0.222 0.317 0.188 0.273
Traffic 0.673 0.461 0.576 0.396 0.546 0.384 - - 0.610 0.399 0.588 0.412 0.613 0.407 0.615 0.421 - - 0.564 0.386 - -
Solar 0.203 0.278 0.184 0.277 0.168 0.272 0.221 0.252 0.219 0.305 0.505 0.549 0.771 0.604 0.393 0.319 0.411 0.428 0.815 0.710 0.714 0.704
PEMS08 0.324 0.384 0.312 0.374 0.293 0.351 - - 0.812 0.692 1.369 0.979 0.866 0.695 1.707 1.024 - - 1.730 1.066 - -
Wind 1.144 0.765 1.120 0.758 1.042 0.733 1.186 0.772 1.292 0.836 1.251 0.820 1.201 0.783 1.478 0.834 - - 1.337 0.829 1.299 0.795
NYSE 0.527 0.553 0.442 0.430 0.529 0.507 0.880 0.642 0.620 0.626 - - 0.988 0.704 1.129 0.720 - - - - 0.492 0.441
$1^{st}$ Count 8 8 6 4 15 9 9 9 0 3 3 2 0 0 0 0 1 2 2 3 0 4
$2^{nd}$ Count 3 1 15 2 10 9 2 7 8 6 2 3 0 0 0 0 2 1 5 3 4 4
Table 2: Average results of zero-shot probabilistic forecasting experiments on datasets from ProbTS. Lower CRPS or NMAE values indicate better predictions. (’-’) denotes datasets included in the model’s pretraining and therefore excluded from testing. (’/’) denotes the excessive time consumption. Red: the best, Blue: the 2nd best. Full results are in Table 14 of Appendix C.
Models FLAME FLAME FLAME Sundial Chronos MOIRAI Lag-Llama TSDiff CSDI TimeGrad GRU NVP
Small Base Large (2025) (2024) (2024) (2023) (2023) (2022) (2022) (2021)
Metrics CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE
ETT (Avg) 0.235 0.293 0.250 0.317 0.236 0.295 0.231 0.273 0.290 0.316 0.366 0.377 0.273 0.310 0.370 0.465 0.304 0.389 0.493 0.619 0.476 0.605
Weather 0.089 0.114 0.087 0.105 0.074 0.090 0.087 0.102 0.142 0.158 0.179 0.143 0.096 0.106 0.132 0.134 0.077 0.093 0.125 0.155 0.119 0.147
Electricity 0.093 0.118 0.086 0.110 0.081 0.097 0.081 0.098 - - 0.247 0.290 - - 0.407 0.519 / / 0.102 0.126 0.101 0.127
Traffic 0.268 0.337 0.248 0.306 0.247 0.308 - - 0.269 0.295 - - 0.330 0.385 0.327 0.392 / / 0.211 0.246 0.198 0.245
Exchange 0.043 0.048 0.040 0.047 0.042 0.047 0.045 0.049 0.044 0.047 0.045 0.050 0.057 0.069 0.084 0.111 0.069 0.086 0.082 0.095 0.073 0.093
ILI 0.162 0.198 0.130 0.166 0.154 0.188 0.148 0.166 0.170 0.197 0.159 0.197 0.156 0.211 0.248 0.259 0.276 0.290 0.284 0.310 0.283 0.309
$1^{st}$ Count 3 3 7 3 15 14 5 8 0 1 0 1 0 0 0 0 2 2 0 2 4 2
$2^{nd}$ Count 7 5 5 5 7 6 5 8 0 2 1 1 1 0 0 0 6 5 4 2 0 2
4.1 Deterministic Forecasting
As shown in Table 1, the FLAME family consistently outperforms other advanced time series foundation models on datasets from TSFM-Bench. Compared with the recent state-of-the-art models Sundial, FLAME Large outperforms it in most settings, achieving most 1st counts in zero-shot deterministic forecasting tasks. Note that even the FLAME Small (2M) can achieve competitive performance with Sundial Base (128M) while only possessing the 1/64 parameter scale. Compared with the heaviest model Time-MoE Ultra (2.4B), FLAME Small possesses the 1/1,200 parameter scale but achieves average MSE reduction of 15.0%. Considering lightweight baselines like UniTS and ROSE, FLAME Small also shows excellent advantages in generalization capabilities, achieving significant average MSE reductions of 26.3% and 17.4%.
4.2 Probabilistic Forecasting
Table 2 represents the results of probabilistic forecasting on ProbTS. Among time series foundation models, only FLAME and Sundial adopt generative probabilistic modeling and show excellent advantages over other models. Considering the accuracy, FLAME Large achieves the most 1st counts in zero-shot probabilistic forecasting tasks, exceling both the periodic scenarios like ETT and Weather, and unstationary scenarios like Exchange and ILI. Compared with the previous well-recognized baseline MOIRAI Large and Chronos Large, FLAME Large achieves significant average CRPS reductions of 34.3% and 18.1%, demonstrating the strong capability of modeling intricate distributions over the forecasting horizons. We observe that FLAME’s zero-shot performance can even outperform the full-shot baselines a lot. Compared with diffusion-based CSDI and Flow-based GRU NVP, which need dozens of hours to train from scratch on a large dataset, FLAME Large separately achieves average 23.5% and 35.1% CRPS reduction. This demonstrates the effectiveness of Flow-based head in FLAME, which provides both efficiency and accuracy.
4.3 Model Analysis
To reveal where the generalization capability of FLAME comes, we analyze the key components of FLAME including Local-Perception, SSD-Decoder, and Flow-based Head in Table 3. Results show that all of our designs are the most reasonable, see–Appendix B for full results and analytics.
Table 3: Ablation studies on Local-Perception (line 3-5), SSD-Decoder (line 6-8), and Flow-based Head (line 9-11). We choose ETT (Avg), Weather, Electricity and Traffic (shared in TSFM-Bench and ProbTS), and report average MSE and CRPS to evaluate deterministic and probabilisic forecasting.
Variants ETT (Avg) Weather Electricity Traffic
Metrics MSE CRPS MSE CRPS MSE CRPS MSE CRPS
w/o Local-Perception 0.374 0.266 0.278 0.104 0.293 0.109 0.690 0.283
w/o Period Alignment 0.368 0.251 0.248 0.083 0.275 0.096 0.627 0.264
Learnable LegT 0.348 0.245 0.237 0.074 0.222 0.084 0.566 0.253
Self-Attention Decoder 0.391 0.300 0.253 0.106 0.300 0.131 0.654 0.341
Causal-Attention Decoder 0.369 0.260 0.255 0.103 0.230 0.095 0.560 0.266
Mamba v1 Decoder 0.362 0.254 0.256 0.098 0.218 0.087 0.557 0.277
VAE Head 0.366 0.262 0.264 0.093 0.291 0.138 0.745 0.506
Diffusion Head 0.360 0.252 0.230 0.077 0.235 0.088 0.568 0.266
Flow Match Head 0.358 0.251 0.233 0.080 0.232 0.088 0.595 0.297
FLAME Large 0.344 0.236 0.227 0.074 0.217 0.081 0.546 0.247
4.4 Scability Analysis
Though the FLAME family possesses extremely lightweight backbones while achieving state-of-the-art performance, we reveal that it also possesses the potential of scability. As shown in Figure 4, the training curves of the FLAME family indicate that the model capacity increases with the parameter scale. According to the results in Table 12, the zero-shot performance also gets better and the FLAME Large achieves most 1st Counts in both deterministic and probabilistic forecasting tasks.
Figure 4: The training curves of the FLAME family.
Since FLAME models distributions over the forecasting horizon, we also study the scability of generative forecasting by exploring the correlations between the sampling number and performance–see Figure 5. The results show that the CPRS and NMAE keep consistent promotion as the sampling number increases, and achieve good performance in the number of 100, which may not cause large efficiency burdens in real-world applications.
Figure 5: We show the CRPS and NMAE on ProbTS w.r.t. the sampling number. Full results are in Table 9 of Appendix B.
4.5 Efficiency Analysis
As a characteristic property of FLAME, the extreme efficiency owes to the lightweight backbone and Flow-based Head. For detailed quantification, we report the Parameter scale, Multiply-Accumulate Operations (MACs), Max GPU Memory Occupation, and Inference Time in Table 4. We compare the FLAME Large with time series foundation models and end-to-end supervised probabilistic models.
Results show FLAME Large achieves consistent efficiency advantages over other time series foundation models. Compared with end-to-end supervised models like TimeGrad and CSDI, FLAME Large also has less parameters, faster inference speed, and less MACs and GPU Memory.
Table 4: Efficiency analysis of FLAME and other baselines on ETTm1 dataset, evaluated with the horizon of 720 and batch size of 1. We report the Parameter scale, MACs, Max GPU Memory, and Inference Time.
Models Parameters MACs GPU (MB) Inference (s)
Timer 67.4 M 52.6 G 1,435 0.08
TimesFM 200 M 624.9 G 1,395 0.16
Chronos Large 700 M 92327.9 G 10,269 34.33
MOIRAI Large 300 M 97.36 G 2,009 0.1
Time-MoE Ultra 453 M 5252.9 G 14,131 2.13
Sundial Base 128 M 432.3 G 603 0.1
TimeGrad 12.3 M 19.3 G 522 28.61
CSDI 15.9 M 37.5 G 817 86.18
FLAME Large 10.6 M 0.238 G 129 0.037
5 Related works
5.1 Time Series Forecasting
Deterministic Time Series Forecasting is critical in decision intelligence, which has facilitated a surge of modern deep learning methods like PatchTST (Nie et al., 2023), Pathformer (Chen et al., 2024b), and DUET (Qiu et al., 2025e). LMU (Voelker et al., 2019) and FiLM (Zhou et al., 2022) first introduce Legendre Memory (LegT) in time series forecasting but they consider compressing the whole input series into state vectors, which may cause the information loss. Compared with them, FLAME leverages the LegT in compressing varying token-wise environmental information and aligning them with the original tokens, thus adapting to the multiple frequencies in pretraining corpus and obtaining strong generalization capabilities. Another important direction is Probabilisic Time Series Forecasting, which can quantify the uncertainties and produce more practical significance in real-world applications. Recently, generative mechanims such as VAE, Diffusion, and Flow are widely studied. Among them, $D^{3}$VAE (Li et al., 2022) and $K^{2}$VAE (Wu et al., 2025b) are the most advanced VAE models in probabilistic time series forecasting. Considering multi-step generative mechanisms like Diffusion and Normalization Flow, recent models include GRU NVP (Rasul et al., 2021b), CSDI (Tashiro et al., 2021), TimeGrad (Rasul et al., 2021a), and TSDiff (Kollovieh et al., 2023), showing strong performance. Compared with them, FLAME is the first foundation model to adapt normalization flow for token-wise probabilistic modeling.
5.2 Time Series Foundation Models
Recent studies on Time Series Foundation Models manage to reveal the source of generalization capability. Distinct Tranformer-based architectures are devised and heavily stacked, and are trained on billion- or trillion-scale time series corpus, to make efforts in capturing the inherent inductive bias within data. Among them, TimesFM (Das et al., 2024), Timer (Liu et al., 2024c), and Time-MoE (Shi et al., 2024) are strong deterministic forecasting baselines, with hevay backbones and strong zero-shot performance. However, works like UniTS (Gao et al., 2024), TinyTTM (Ekambaram et al., 2024), ROSE (Wang et al., 2025c) also demonstrate that light backbones can also achieve state-of-the-art deterministic forecasting performance if the inductive bias is captured in a more efficient way. Some time series foundation models support probabilistic forecasting, among them, Lag-Llama (Rasul et al., 2023), Chronos (Ansari et al., ), Moirai (Woo et al., 2024), and Sundial (Liu et al., 2025d) have outperformed some end-to-end supervised probabilistic models. However, all of them possess hevay backbones with hundreds-million to billion parameters, which are not very efficient in making generative probabilistic forecasts on large downstream datasets. Moreover, no existing methods explore how to achieve outstanding probabilistic forecasting performance with lightweight architectures. In this paper, we propose FLAME to handle this bottleneck.
6 Conclusion
In this work, we propose a family of extremely lightweight and highly capable time series foundation models, named FLAME. To sum up, FLAME adapts the Legendre Memory in the backbone to compress varying local environmental information into time series tokens and support adaptive memory utilization in long-range forecasting, thus obtaining strong generalization capabilities. FLAME also supports probabilistic time series forecasting through a Normalization Flow based forecasting head. Comprehensive experiments on deterministic and probabilistic forecasting tasks demonstrate that FLAMEs are strong out-of-the-box tools of decision intelligence, possessing both accuracy and efficiency.
In the future work, we will further explore the non-deterministic modeling in all the tasks of time series analysis, e.g., probabilistic imputation, time series generation, anomaly detection, and classification. Another direction is to develop more lightweight and robust mechanisms to support generative modeling, and extend them via pretraining to the out-of-the-box paradigm.
Appendix A Implementation Details
A.1 Pretraining Corpus
To pre-train our model, we utilize a broad collection of time series datasets drawn from multiple sources, including selected subsets from the Monash (Godahewa et al., 2021b), UEA (Bagnall et al., 2018), and UCR (Dau et al., 2019) repositories, along with several widely-used benchmarks (Zhang et al., 2017; Wang et al., 2024; Liu et al., 2022; McCracken & Ng, 2016; Taieb et al., 2012). A full inventory of these datasets is provided in Table 5. We ensure that none of the pre-training datasets overlap with those used in downstream evaluations. It’s also worth noting that although both the pre-training and target sets include weather data, the former is univariate while the latter contains multivariate sequences. We categorize the pre-training datasets into six domain-specific groups: Energy, Nature, Health, Transport, and Web. These datasets vary significantly in sampling rates, ranging from milliseconds to monthly intervals, capturing the diversity and richness of real-world time series. All datasets are decomposed into univariate sequences and FLAME is trained using a Channel Independent (CI) strategy.
Table 5: List of pretraining datasets.
Domain Dataset Frequency Piont Number Source
Energy Aus. Electricity Demand Half Hourly 1155264 Monash (Godahewa et al., 2021b)
Wind 4 Seconds 7397147 Monash (Godahewa et al., 2021b)
Wind Farms Minutely 172178060 Monash (Godahewa et al., 2021b)
Solar Power 4 Seconds 7397222 Monash (Godahewa et al., 2021b)
London Smart Meters Half Hourly 166527216 Monash (Godahewa et al., 2021b)
Economic FRED_MD Monthly 77896  (McCracken & Ng, 2016)
Bitcoin Daily 75364 Monash (Godahewa et al., 2021b)
NN5 Daily 87801  (Taieb et al., 2012)
Health MotorImagery 0.001 Seconds 72576000 UEA (Bagnall et al., 2018)
SelfRegulationSCP1 0.004 Seconds 3015936 UEA (Bagnall et al., 2018)
SelfRegulationSCP2 0.004 Seconds 3064320 UEA (Bagnall et al., 2018)
AtrialFibrillation 0.008 Seconds 38400 UEA (Bagnall et al., 2018)
PigArtPressure - 624000 UCR (Dau et al., 2019)
PIGCVP - 624000 UCR (Dau et al., 2019)
TDbrain 0.002 Seconds 79232703  (Wang et al., 2024)
Transport Pems03 5 Minute 9382464  (Liu et al., 2022)
Pems07 5 Minute 24921792  (Liu et al., 2022)
Pems-bay 5 Minute 16937700  (Liu et al., 2022)
Pedestrian_Counts Hourly 3132346 Monash (Godahewa et al., 2021b)
Web Web Traffic Daily 116485589 Monash (Godahewa et al., 2021b)
Nature Phoneme - 2160640 UCR(Dau et al., 2019)
EigenWorms - 27947136 UEA (Bagnall et al., 2018)
PRSA Hourly 4628448  (Zhang et al., 2017)
Temperature Rain Daily 23252200 Monash (Godahewa et al., 2021b)
StarLightCurves - 9457664 UCR (Dau et al., 2019)
Worms 0.033 Seconds 232200 UCR (Dau et al., 2019)
Saugeen River Flow Daily 23741 Monash (Godahewa et al., 2021b)
Sunspot Daily 73924 Monash (Godahewa et al., 2021b)
Weather Daily 43032000 Monash (Godahewa et al., 2021b)
KDD Cup 2018 Daily 2942364 Monash(Godahewa et al., 2021b)
US Births Daily 7305 Monash (Godahewa et al., 2021b)
A.2 Baselines
For zero-shot deterministic forecasting, we compare the FLAME family against 8 advanced foundation models: Sundial (Liu et al., 2025d), ROSE (Wang et al., 2025c), Timer (Liu et al., 2024c), MOIRAI (Woo et al., 2024), Chronos (Ansari et al., ), LightGTS (Wang et al., 2025b), Time-MoE (Shi et al., 2024), and TinyTTM (Ekambaram et al., 2024). For zero-shot probabilistic forecasting, we compare the FLAME family with 4 foundation models: Sundial (Liu et al., 2025d), MOIRAI (Woo et al., 2024), Chronos (Ansari et al., ), and Lag-Llama (Rasul et al., 2023), and 4 advanced end-to-end supervised models including TSDiff (Kollovieh et al., 2023), TimeGrad (Rasul et al., 2021a), CSDI (Tashiro et al., 2021), and GRU NVP (Rasul et al., 2021b). The corresponding codebases and implementation details are summarized in Table 6.
Table 6: Code repositories for baselines.
Model Types Models Code Repositories
End2End TSDiff https://github.com/amazon-science/unconditional-time-series-diffusion
CSDI https://github.com/ermongroup/CSDI
TimeGrad https://github.com/Zjh152/TimeGrad
GRU NVP https://github.com/microsoft/ProbTS
Foundation Sundial https://github.com/thuml/Sundial
ROSE https://github.com/decisionintelligence/TSFM-Bench
Timer https://github.com/thuml/Large-Time-Series-Model
MOIRAI https://github.com/redoules/moirai
Chronos https://github.com/amazon-science/chronos-forecasting
LightGTS https://github.com/decisionintelligence/LightGTS
Time-MoE https://github.com/Time-MoE/Time-MoE
Lag-Llama https://github.com/time-series-foundation-models/lag-llama
TinyTTM https://huggingface.co/ibm-granite/granite-timeseries-ttm-r1
A.3 Benchmarks
To thoroughly assess the effectiveness of FLAME, we conduct experiments on TSFM-Bench (Li et al., 2025c) and ProbTS (Zhang et al., 2024). Our evaluation spans 11 deterministic and 9 probabilistic forecasting datasets.
For deterministic forecasting, we use ETTm1, ETTm2, ETTh1, ETTh2, Weather, Electricity, Traffic, Solar, PEMS08, Wind, and NYSE from TSFM-Bench. For most datasets, the prediction length is set to $L\in\{96,192,336,720\}$, while NYSE uses $L\in\{24,36,48,60\}$.
For probabilistic forecasting, we adopt ETTm1, ETTm2, ETTh1, ETTh2, Weather, Electricity, Traffic, Exchange, and ILI from ProbTS. The prediction length is set to $L\in\{96,192,336,720\}$ for most datasets, while ILI-L uses $L\in\{24,36,48,60\}$.
All models are configured with the contextual length of the best performance recommended in their papers. It is important to clarify that datasets shared in both the TSFM-Bench and ProbTS (e.g., ETTh1, Traffic) are different in evaluation settings like strides. A summary of dataset statistics can be found in Table 7.
Table 7: Statistics of benchmark datasets.
Dataset Domain Frequency Lengths Dim Split Stride Benchmark Description
ETTm1 Electricity 15 mins 57,600 7 6:2:2 1 TSFM-Bench Power transformer 1, comprising seven indicators such as oil temperature and useful load
ETTm2 Electricity 15 mins 57,600 7 6:2:2 1 TSFM-Bench Power transformer 2, comprising seven indicators such as oil temperature and useful load
ETTh1 Electricity 1 hour 14,400 7 6:2:2 1 TSFM-Bench Power transformer 1, comprising seven indicators such as oil temperature and useful load
ETTh2 Electricity 1 hour 14,400 7 6:2:2 1 TSFM-Bench Power transformer 2, comprising seven indicators such as oil temperature and useful load
Weather Environment 10 mins 52,696 21 7:1:2 1 TSFM-Bench Recorded every for the whole year 2020, which contains 21 meteorological indicators
Electricity Electricity 1 hour 26,304 321 7:1:2 1 TSFM-Bench Electricity records the electricity consumption in kWh every 1 hour from 2012 to 2014
Traffic Traffic 1 hour 17,544 862 7:1:2 1 TSFM-Bench Road occupancy rates measured by 862 sensors on San Francisco Bay area freeways
Solar Energy 10 mins 52,560 137 6:2:2 1 TSFM-Bench Solar production records collected from 137 PV plants in Alabama
PEMS08 Traffic 5 mins 17,856 170 6:2:2 1 TSFM-Bench Traffic flow time series collected from the CalTrans PeMS
Wind Energy 15 mins 48,673 7 7:1:2 1 TSFM-Bench Wind power records from 2020-2021 at 15-minute intervals
NYSE Stock 1 day 1,243 5 7:1:2 1 TSFM-Bench Records opening price, closing price, trading volume, lowest price, and highest price
ETTm1 Electricity 15 mins 57,600 7 6:2:2 96 ProbTS Power transformer 1, comprising seven indicators such as oil temperature and useful load
ETTm2 Electricity 15 mins 57,600 7 6:2:2 96 ProbTS Power transformer 2, comprising seven indicators such as oil temperature and useful load
ETTh1 Electricity 1 hour 14,400 7 6:2:2 96 ProbTS Power transformer 1, comprising seven indicators such as oil temperature and useful load
ETTh2 Electricity 1 hour 14,400 7 6:2:2 96 ProbTS Power transformer 2, comprising seven indicators such as oil temperature and useful load
Weather Environment 10 mins 52,696 21 7:1:2 96 ProbTS Recorded every for the whole year 2020, which contains 21 meteorological indicators
Electricity Electricity 1 hour 26,304 321 7:1:2 96 ProbTS Electricity records the electricity consumption in kWh every 1 hour from 2012 to 2014
Traffic Traffic 1 hour 17,544 862 7:1:2 96 ProbTS Road occupancy rates measured by 862 sensors on San Francisco Bay area freeways
Exchange Economic 1 day 7,588 8 7:1:2 96 ProbTS ExchangeRate collects the daily exchange rates of eight countries
ILI Health 1 week 966 7 7:1:2 96 ProbTS Recorded indicators of patients data from Centers for Disease Control and Prevention
A.4 Experimental Settings
Pretraining
We implement the FLAME family using Distributed Data Parallel in PyTorch (Paszke et al., 2019), and conduct all experiments on 8 NVIDIA A800 GPUs with 80GB GPU memory. The model is optimized using the AdamW optimizer with an initial learning rate of $5\times 10^{-5}$. To progressively reduce the learning rate during training, we employ a step decay schedule via the StepLR scheduler. During pre-training, we use 12 historical tokens and 4 prediction tokens, with a reference patch size $p=48$. The batch size is set to 8,192. Full details on model configurations and parameter sizes are summarized in Table 8.
Table 8: Detailed model configurations of the FLAME family and corresponding parameter counts.
Models Encoder Layers Decoder Layers Couple Layers Model Dim $d$ FFN Dim $d^{\prime}$ Parameters
FLAME Small 1 1 3 256 512 2M
FLAME Base 1 3 5 256 512 6M
FLAME Large 1 6 7 256 512 10M
Downstream Forecasting
For downstream forecasting tasks, we apply periodic patching strategies adapted to each dataset’s temporal characteristics. The number of past tokens remains fixed at 11, while the prediction lengths are evaluated at 96, 192, 336, and 720 steps.
We also address the "Drop Last" issue highlighted in recent studies (Qiu et al., 2024, 2025c; Li et al., 2025c), where setting drop_last=True during test evaluation may lead to misleading results due to incomplete batches. To ensure consistency and fairness, we set drop_last=False for all baseline models in our experiments.
A.5 Evaluation Metrics
Regarding evaluation metrics, following the experimental setup in TSFM-Bench, we adopt Mean Squared Error (MSE) and Mean Absolute Error (MAE) as evaluation metrics for deterministic forecasting. For probabilistic forecasting, we use Continuous Ranked Probability Score (CRPS) and Normalized Mean Absolute Error (NMAE) in ProbTS. Considering the case with $K$ variates and $T$ forecasting horizon.
Mean Squared Error (MSE)
The Mean Squared Error quantifies the average of the squared differences between predicted values and their corresponding ground truths. It penalizes larger errors more heavily due to the squaring operation, making it sensitive to outliers. Formally, it is defined as:
$\textrm{MSE}=\frac{1}{K\times T}\sum_{k=1}^{K}\sum_{t=1}^{T}(x^{k}_{t}-\hat{x}^{k}_{t})^{2},$ (33)
where $K$ denotes the number of variables, $T$ the prediction horizon, $x^{k}_{t}$ the true value, and $\hat{x}^{k}_{t}$ the predicted value.
Mean Absolute Error (MAE)
The Mean Absolute Error calculates the average magnitude of prediction errors without considering their direction. By focusing on the absolute differences, MAE provides a robust and interpretable measure of accuracy:
$\textrm{MAE}=\frac{1}{K\times T}\sum_{k=1}^{K}\sum_{t=1}^{T}|x^{k}_{t}-\hat{x}^{k}_{t}|,$ (34)
where all terms follow the same definition as above. Unlike MSE, MAE treats all errors equally and is less sensitive to large deviations.
Continuous Ranked Probability Score (CRPS)
The CRPS assesses the quality of probabilistic forecasts by comparing the predicted cumulative distribution function (CDF) $F$ with the observed outcome $x$. It is computed as:
$\textrm{CRPS}=\int_{\mathds{R}}(F(z)-\mathds{I}\{x\leq z\})^{2}dz,$ (35)
where $\mathds{I}\{x\leq z\}$ is the indicator function. CRPS rewards distributions that assign high probability to the true value and achieves its minimum when the predicted distribution matches the true distribution. In practice, we approximate CRPS using the empirical CDF $\hat{F}(z)=\frac{1}{n}\sum_{i=1}^{n}\mathds{I}\{X_{i}\leq z\}$ based on $n=100$ samples drawn from the conditional predictive distribution $p_{\theta}(\bm{x}_{t}|\bm{h}_{t})$.
Normalized Mean Absolute Error (NMAE)
The NMAE extends MAE by normalizing it against the total magnitude of ground-truth values, thereby enabling fair comparison across datasets with varying scales. Its formula is:
$\textrm{NMAE}=\frac{\sum_{k=1}^{K}\sum_{t=1}^{T}|x^{k}_{t}-\hat{x}^{k}_{t}|}{\sum_{k=1}^{K}\sum_{t=1}^{T}|x^{k}_{t}|}$ (36)
Appendix B Full Model Analytics
B.1 Analytics of Sampling Number
Given that FLAME models distributions across the forecasting horizon, we conduct an in-depth study on the scalability of generative forecasting by examining the correlations between the sampling number and performance, as presented in Table 9. Specifically, we choose four sampling numbers: 10, 20, 50, and 100. The findings indicate that both CPRS and NMAE demonstrate a consistent improvement as the sampling number rises. Moreover, a sampling number of 100 leads to excellent performance, and this is unlikely to impose significant efficiency burdens in practical applications.
Table 9: Full results of probabilistic forecasting experiments with different sampling numbers.
Models FLAME-10 FLAME-20 FLAME-50 FLAME-100
Metric CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE
96 1.255 0.412 0.388 0.477 0.323 0.418 0.315 0.407
192 0.905 0.444 0.372 0.433 0.358 0.432 0.337 0.426
336 0.369 0.468 0.448 0.682 0.352 0.448 0.351 0.447
720 0.388 0.486 0.394 0.491 0.394 0.488 0.383 0.476
ETTm1
avg 0.729 0.453 0.401 0.521 0.357 0.447 0.347 0.439
96 0.134 0.173 0.129 0.167 0.125 0.163 0.124 0.155
192 0.208 0.196 0.187 0.190 0.144 0.186 0.143 0.175
336 0.477 0.212 0.427 0.208 0.159 0.204 0.158 0.193
720 0.662 0.236 0.631 0.232 0.187 0.229 0.184 0.219
ETTm2
avg 0.370 0.204 0.344 0.199 0.154 0.196 0.152 0.186
96 0.299 0.346 0.269 0.331 0.262 0.324 0.259 0.310
192 0.294 0.365 0.285 0.354 0.279 0.347 0.278 0.342
336 0.309 0.383 0.298 0.372 0.293 0.365 0.290 0.362
720 0.328 0.406 0.317 0.396 0.311 0.390 0.303 0.367
ETTh1
avg 0.308 0.375 0.292 0.363 0.286 0.357 0.283 0.345
96 0.158 0.200 0.150 0.191 0.146 0.188 0.144 0.186
192 0.167 0.211 0.162 0.207 0.160 0.206 0.159 0.204
336 0.181 0.226 0.176 0.222 0.174 0.221 0.172 0.217
720 0.191 0.234 0.187 0.231 0.183 0.229 0.182 0.228
ETTh2
avg 0.174 0.218 0.169 0.213 0.166 0.211 0.164 0.209
96 0.080 0.098 0.178 0.093 0.073 0.091 0.067 0.083
192 0.103 0.112 0.092 0.105 0.084 0.103 0.068 0.082
336 1.382 0.121 0.233 0.117 0.092 0.114 0.079 0.094
720 2.471 0.133 1.384 0.129 0.101 0.127 0.083 0.100
Weather
avg 1.009 0.116 0.472 0.111 0.088 0.109 0.074 0.090
96 1.734 0.097 0.072 0.092 0.070 0.090 0.069 0.089
192 0.093 0.119 1.662 0.116 0.086 0.112 0.072 0.088
336 0.662 0.147 0.375 0.143 0.107 0.140 0.081 0.096
720 2.406 0.173 1.354 0.168 0.125 0.166 0.100 0.116
Electricity
avg 1.224 0.134 0.866 0.130 0.097 0.127 0.081 0.097
96 0.562 0.301 0.491 0.309 0.224 0.273 0.220 0.268
192 0.811 0.324 0.681 0.333 0.240 0.298 0.238 0.294
336 2.911 0.348 3.284 0.364 0.257 0.323 0.254 0.319
720 3.491 0.378 2.824 0.484 0.278 0.354 0.276 0.351
Traffic
avg 1.944 0.338 1.820 0.373 0.250 0.312 0.247 0.308
96 0.023 0.027 0.022 0.027 0.022 0.027 0.020 0.026
192 0.032 0.037 0.032 0.037 0.031 0.037 0.031 0.035
336 0.045 0.049 0.044 0.049 0.044 0.049 0.044 0.046
720 1.234 0.098 0.145 0.082 0.074 0.079 0.073 0.079
Exchange
avg 0.334 0.053 0.061 0.049 0.043 0.048 0.042 0.047
24 0.156 0.188 0.157 0.197 0.157 0.194 0.136 0.168
36 0.178 0.216 0.181 0.225 0.177 0.214 0.145 0.179
48 0.197 0.234 0.201 0.244 0.196 0.235 0.164 0.200
60 0.196 0.234 0.190 0.231 0.192 0.233 0.169 0.204
ILI
avg 0.182 0.218 0.182 0.224 0.181 0.219 0.154 0.188
B.2 Analytics of Local-Perception
As shown in Table 10, we select three variants to assess the Local-Perception. These variants include eliminating the entire Local-Perception (termed as “w/o Local-Perception”), encoding the closest single period without performing alignment (referred to as “w/o Period Alignment”), and making the initialized matrices $A,B$ learnable (denoted as “Learnable LegT”). Our observations indicate that Local-Perception indeed has a significant impact. Specifically, the variant lacking Local-Perception experiences a collapse in zero-shot performance. Period Alignment is also of importance. When considering datasets such as Electricity and Traffic, which contain complex periodic information, the zero-shot performance deteriorates substantially without the assistance of Period Alignment. Regarding Learnable LegT, its performance shows a slight decline when the initialized matrices $A,B$ are made learnable. This is because, upon being pretrained on a large-scale time series corpus, the Legendre coefficients within the matrices may gradually change, thereby losing the on-line approximation capability.
Table 10: Experimental results on Local Perception analysis, comparing configurations without Local Perception (LP), without Period Alignment (P.A.), and with Learnable LegT ($L.$).
Models w/o LP w/o P.A. with L.L. FLAME Large
Metric MSE CRPS MSE CRPS MSE CRPS MSE CRPS
96 0.365 0.352 0.362 0.332 0.297 0.317 0.301 0.315
192 0.389 0.385 0.387 0.364 0.327 0.335 0.328 0.337
336 0.417 0.407 0.408 0.398 0.353 0.359 0.351 0.351
720 0.470 0.462 0.483 0.423 0.397 0.390 0.392 0.383
ETTm1
avg 0.410 0.402 0.410 0.379 0.344 0.350 0.343 0.347
96 0.229 0.144 0.216 0.133 0.178 0.121 0.186 0.124
192 0.283 0.154 0.258 0.149 0.253 0.146 0.247 0.143
336 0.311 0.186 0.304 0.171 0.297 0.164 0.300 0.158
720 0.362 0.200 0.363 0.195 0.365 0.191 0.357 0.184
ETTm2
avg 0.296 0.171 0.285 0.162 0.273 0.156 0.273 0.152
96 0.372 0.283 0.362 0.279 0.352 0.255 0.358 0.259
192 0.418 0.304 0.402 0.291 0.408 0.291 0.392 0.278
336 0.433 0.322 0.428 0.300 0.413 0.304 0.411 0.290
720 0.450 0.376 0.459 0.318 0.438 0.328 0.444 0.303
ETTh1
avg 0.418 0.321 0.413 0.297 0.403 0.295 0.401 0.283
96 0.312 0.148 0.303 0.145 0.311 0.161 0.300 0.144
192 0.368 0.167 0.358 0.159 0.364 0.178 0.351 0.159
336 0.384 0.182 0.375 0.174 0.391 0.199 0.374 0.172
720 0.425 0.188 0.422 0.186 0.422 0.185 0.416 0.182
ETTh2
avg 0.372 0.171 0.365 0.166 0.372 0.181 0.360 0.164
96 0.233 0.076 0.195 0.071 0.168 0.067 0.152 0.067
192 0.270 0.099 0.225 0.074 0.214 0.069 0.199 0.068
336 0.295 0.112 0.261 0.088 0.255 0.078 0.250 0.079
720 0.312 0.127 0.311 0.098 0.312 0.081 0.307 0.083
Weather
avg 0.278 0.104 0.248 0.083 0.237 0.074 0.227 0.074
96 0.214 0.088 0.202 0.077 0.166 0.072 0.161 0.069
192 0.253 0.098 0.237 0.091 0.198 0.073 0.190 0.072
336 0.309 0.112 0.288 0.104 0.233 0.085 0.227 0.081
720 0.397 0.138 0.374 0.112 0.292 0.105 0.289 0.100
Electricity
avg 0.293 0.109 0.275 0.096 0.222 0.084 0.217 0.081
96 0.599 0.239 0.557 0.229 0.503 0.222 0.480 0.220
192 0.631 0.275 0.592 0.251 0.526 0.246 0.511 0.238
336 0.722 0.301 0.647 0.282 0.589 0.265 0.553 0.254
720 0.808 0.318 0.712 0.294 0.647 0.279 0.639 0.276
Traffic
avg 0.690 0.283 0.627 0.264 0.566 0.253 0.546 0.247
B.3 Analytics of SSD-Decoder
Subsequently, we conduct a comparison between the SSD-Decoder and the decoders of other mechanisms in Table 11. More precisely, we select the Self-Attention Decoder, Causal-Attention Decoder, and Mamba v1 Decoder as the baselines. In the context of forecasting tasks, causal correlations are more suitable. As a result, the Self-Attention Decoder exhibits the poorest performance. We further note that the SSD-Decoder outperforms both the Mamba v1 Decoder and the Causal-Attention Decoder, which showcases its superior ability to leverage memory during the forecasting process.
Table 11: Forecasting performance comparison of different decoders: Self-Attention (SA), Causal-Attention (CA), Mamba v1, and our SSD-Decoder used in FLAME.
Models SA CA Mamba v1 FLAME
Metric MSE CRPS MSE CRPS MSE CRPS MSE CRPS
96 0.389 0.448 0.377 0.358 0.337 0.332 0.301 0.315
192 0.421 0.467 0.395 0.382 0.359 0.341 0.328 0.337
336 0.449 0.485 0.413 0.401 0.390 0.367 0.351 0.351
720 0.491 0.508 0.461 0.442 0.444 0.390 0.392 0.383
ETTm1
avg 0.438 0.477 0.412 0.396 0.383 0.358 0.343 0.347
96 0.235 0.167 0.209 0.141 0.212 0.139 0.186 0.124
192 0.272 0.198 0.251 0.167 0.256 0.162 0.247 0.143
336 0.352 0.223 0.313 0.182 0.292 0.179 0.300 0.158
720 0.398 0.257 0.368 0.224 0.356 0.198 0.357 0.184
ETTm2
avg 0.314 0.211 0.285 0.179 0.279 0.170 0.273 0.152
96 0.358 0.269 0.354 0.263 0.363 0.271 0.358 0.259
192 0.404 0.302 0.395 0.286 0.404 0.303 0.392 0.278
336 0.420 0.331 0.423 0.297 0.428 0.325 0.411 0.290
720 0.455 0.352 0.475 0.325 0.468 0.344 0.444 0.303
ETTh1
avg 0.409 0.314 0.412 0.293 0.416 0.311 0.401 0.283
96 0.336 0.175 0.288 0.139 0.303 0.148 0.300 0.144
192 0.382 0.203 0.362 0.170 0.370 0.172 0.351 0.159
336 0.414 0.212 0.389 0.177 0.401 0.186 0.374 0.172
720 0.472 0.209 0.432 0.198 0.409 0.202 0.416 0.182
ETTh2
avg 0.401 0.200 0.368 0.171 0.371 0.177 0.360 0.164
96 0.188 0.088 0.209 0.093 0.207 0.087 0.152 0.067
192 0.221 0.094 0.237 0.097 0.237 0.091 0.199 0.068
336 0.275 0.113 0.261 0.105 0.269 0.102 0.250 0.079
720 0.328 0.127 0.311 0.117 0.311 0.112 0.307 0.083
Weather
avg 0.253 0.106 0.255 0.103 0.256 0.098 0.227 0.074
96 0.210 0.115 0.172 0.075 0.163 0.072 0.161 0.069
192 0.255 0.127 0.204 0.084 0.191 0.077 0.190 0.072
336 0.318 0.132 0.243 0.098 0.234 0.090 0.227 0.081
720 0.415 0.149 0.302 0.122 0.285 0.108 0.289 0.100
Electricity
avg 0.300 0.131 0.230 0.095 0.218 0.087 0.217 0.081
96 0.587 0.311 0.492 0.238 0.501 0.244 0.480 0.220
192 0.617 0.338 0.521 0.256 0.522 0.270 0.511 0.238
336 0.662 0.343 0.572 0.275 0.562 0.294 0.553 0.254
720 0.749 0.372 0.655 0.294 0.642 0.298 0.639 0.276
Traffic
avg 0.654 0.341 0.560 0.266 0.557 0.277 0.546 0.247
B.4 Analytics of Flow-based Head
To assess the efficacy of the Flow-based Head, we choose alternative generative mechanisms to implement forecasting heads as baselines–see Table 12. These include Variational AutoEncoder (VAE), Diffusion, and Flow Match. Given that the VAE adheres to a one-step generation paradigm, it exhibits subpar performance in scenarios such as Electricity and Traffic. This is because accurately generating distributions over the complex real-world forecasting horizon in these scenarios poses a significant challenge. Regarding Diffusion and Flow Match, although they possess robust multi-step generation paradigms, our findings indicate that they are not as potent as our meticulously-designed Flow-based Head.
Table 12: Forecasting performance comparison of various generative heads: VAE, Diffusion, Flow Match, and our Flow-based Head used in FLAME.
Models VAE Diffusion Flow Match FLAME
Metric MSE CRPS MSE CRPS MSE CRPS MSE CRPS
96 0.334 0.348 0.305 0.318 0.311 0.323 0.301 0.315
192 0.364 0.382 0.340 0.345 0.331 0.340 0.328 0.337
336 0.388 0.397 0.359 0.362 0.361 0.360 0.351 0.351
720 0.421 0.448 0.401 0.390 0.398 0.386 0.392 0.383
ETTm1
avg 0.377 0.394 0.351 0.354 0.350 0.352 0.343 0.347
96 0.223 0.152 0.194 0.128 0.215 0.139 0.186 0.124
192 0.292 0.179 0.261 0.145 0.257 0.144 0.247 0.143
336 0.345 0.197 0.322 0.163 0.309 0.168 0.300 0.158
720 0.392 0.215 0.370 0.200 0.381 0.205 0.357 0.184
ETTm2
avg 0.313 0.186 0.287 0.159 0.291 0.164 0.273 0.152
96 0.366 0.279 0.359 0.261 0.363 0.267 0.358 0.259
192 0.398 0.285 0.394 0.288 0.404 0.295 0.392 0.278
336 0.423 0.301 0.428 0.318 0.427 0.330 0.411 0.290
720 0.457 0.320 0.479 0.348 0.462 0.327 0.444 0.303
ETTh1
avg 0.411 0.296 0.415 0.304 0.414 0.305 0.401 0.283
96 0.296 0.142 0.332 0.169 0.328 0.164 0.300 0.144
192 0.358 0.167 0.378 0.180 0.364 0.168 0.351 0.159
336 0.379 0.182 0.394 0.198 0.384 0.191 0.374 0.172
720 0.425 0.197 0.448 0.215 0.428 0.202 0.416 0.182
ETTh2
avg 0.365 0.172 0.388 0.191 0.376 0.181 0.360 0.164
96 0.176 0.083 0.149 0.068 0.155 0.072 0.152 0.067
192 0.229 0.086 0.199 0.069 0.206 0.074 0.199 0.068
336 0.292 0.098 0.256 0.083 0.257 0.084 0.250 0.079
720 0.357 0.105 0.314 0.089 0.312 0.088 0.307 0.083
Weather
avg 0.264 0.093 0.230 0.077 0.233 0.080 0.227 0.074
96 0.241 0.129 0.172 0.078 0.167 0.075 0.161 0.069
192 0.260 0.134 0.199 0.079 0.195 0.079 0.190 0.072
336 0.304 0.142 0.251 0.088 0.246 0.086 0.227 0.081
720 0.357 0.147 0.318 0.105 0.318 0.110 0.289 0.100
Electricity
avg 0.291 0.138 0.235 0.088 0.232 0.088 0.217 0.081
96 0.721 0.482 0.488 0.236 0.507 0.248 0.480 0.220
192 0.742 0.498 0.528 0.264 0.557 0.296 0.511 0.238
336 0.747 0.512 0.572 0.270 0.603 0.305 0.553 0.254
720 0.769 0.532 0.682 0.293 0.712 0.337 0.639 0.276
Traffic
avg 0.745 0.506 0.568 0.266 0.595 0.297 0.546 0.247
Appendix C Full Results
C.1 Full zero-shot results
We provide all the results of the zero-shot deterministic forecasting in Table 13. As shown in Table 13, we include 11 representative real-world datasets from TSFM-Bench, demonstrating that FLAME achieves state-of-the-art forecasting performance.
Table 13: Full results of zero-shot deterministic forecasting experiments on datasets from TSFM-Bench. Lower MSE or MAE values indicate better predictions. (’-’) denotes datasets included in the model’s pretraining and therefore excluded from testing. Red: the best, Blue: the 2nd best.
Models FLAME FLAME FLAME Sundial LightGTS ROSE Timer Chronos Time-MoE TinyTTM MOIRAI
Small Base Large (2025) (2025) (2025) (2024) (2024) (2024) (2024) (2024)
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTm1
96 0.305 0.353 0.298 0.351 0.301 0.353 0.280 0.334 0.307 0.359 0.512 0.460 0.817 0.611 0.402 0.373 0.281 0.341 0.738 0.541 0.353 0.363
192 0.328 0.370 0.324 0.369 0.328 0.373 0.321 0.366 0.332 0.374 0.512 0.462 0.927 0.659 0.510 0.435 0.305 0.358 0.698 0.547 0.376 0.380
336 0.349 0.384 0.346 0.386 0.351 0.389 0.350 0.389 0.353 0.393 0.523 0.470 1.043 0.704 0.590 0.477 0.369 0.395 0.670 0.533 0.399 0.395
720 0.392 0.416 0.385 0.414 0.392 0.419 0.394 0.418 0.388 0.386 0.552 0.490 1.044 0.722 0.703 0.525 0.469 0.472 0.660 0.550 0.432 0.417
Avg 0.344 0.381 0.338 0.380 0.343 0.384 0.336 0.377 0.345 0.378 0.525 0.471 0.958 0.674 0.551 0.453 0.356 0.392 0.692 0.543 0.390 0.389
ETTm2
96 0.165 0.255 0.171 0.262 0.186 0.272 0.170 0.256 0.166 0.260 0.224 0.309 0.225 0.300 0.192 0.263 0.198 0.288 0.226 0.309 0.189 0.260
192 0.217 0.297 0.227 0.302 0.247 0.315 0.229 0.300 0.222 0.300 0.266 0.333 0.286 0.339 0.256 0.308 0.235 0.312 0.311 0.360 0.247 0.300
336 0.262 0.330 0.280 0.340 0.300 0.352 0.281 0.337 0.267 0.333 0.310 0.358 0.335 0.369 0.315 0.346 0.293 0.348 0.350 0.383 0.295 0.334
720 0.333 0.379 0.348 0.387 0.357 0.394 0.351 0.387 0.340 0.378 0.395 0.407 0.414 0.416 0.409 0.405 0.427 0.428 0.446 0.435 0.372 0.386
Avg 0.244 0.315 0.257 0.323 0.273 0.333 0.258 0.320 0.249 0.318 0.299 0.352 0.315 0.356 0.293 0.331 0.288 0.344 0.333 0.372 0.276 0.320
ETTh1
96 0.362 0.395 0.361 0.392 0.358 0.389 0.348 0.385 0.359 0.390 0.382 0.408 0.454 0.434 0.444 0.409 0.349 0.379 0.364 0.389 0.380 0.398
192 0.406 0.427 0.400 0.422 0.392 0.416 0.393 0.418 0.390 0.409 0.400 0.420 0.522 0.465 0.502 0.443 0.395 0.413 0.386 0.407 0.440 0.434
336 0.436 0.449 0.417 0.436 0.411 0.432 0.422 0.440 0.420 0.436 0.404 0.426 0.559 0.484 0.580 0.460 0.447 0.453 0.404 0.422 0.514 0.474
720 0.471 0.482 0.433 0.459 0.444 0.463 0.481 0.493 0.437 0.460 0.420 0.447 0.714 0.549 0.605 0.495 0.457 0.462 0.424 0.448 0.705 0.568
Avg 0.419 0.438 0.403 0.427 0.401 0.425 0.411 0.434 0.402 0.424 0.402 0.425 0.562 0.483 0.533 0.452 0.412 0.427 0.395 0.417 0.510 0.469
ETTh2
96 0.278 0.339 0.295 0.345 0.300 0.350 0.271 0.333 0.278 0.342 0.298 0.362 0.316 0.359 0.306 0.338 0.292 0.352 0.277 0.335 0.287 0.325
192 0.336 0.381 0.353 0.388 0.351 0.390 0.327 0.376 0.345 0.383 0.336 0.385 0.374 0.398 0.396 0.394 0.347 0.379 0.334 0.373 0.347 0.367
336 0.373 0.410 0.373 0.409 0.374 0.411 0.354 0.402 0.394 0.416 0.353 0.399 0.381 0.410 0.423 0.417 0.406 0.419 0.362 0.402 0.377 0.393
720 0.411 0.441 0.413 0.440 0.416 0.444 0.381 0.435 0.430 0.445 0.395 0.432 0.408 0.434 0.442 0.439 0.439 0.447 0.408 0.444 0.404 0.421
Avg 0.350 0.393 0.359 0.396 0.360 0.399 0.333 0.387 0.362 0.397 0.346 0.395 0.370 0.400 0.392 0.397 0.371 0.399 0.345 0.389 0.354 0.377
Weather
96 0.148 0.204 0.151 0.209 0.152 0.211 0.157 0.205 0.151 0.211 0.200 0.260 0.190 0.236 0.186 0.208 0.157 0.211 0.183 0.242 0.177 0.208
192 0.189 0.243 0.194 0.250 0.199 0.254 0.205 0.251 0.191 0.248 0.239 0.288 0.261 0.293 0.238 0.258 0.208 0.256 0.229 0.285 0.219 0.249
336 0.233 0.278 0.241 0.286 0.250 0.293 0.253 0.289 0.236 0.284 0.279 0.315 0.332 0.340 0.313 0.353 0.255 0.290 0.289 0.330 0.277 0.292
720 0.289 0.320 0.301 0.330 0.307 0.333 0.320 0.336 0.299 0.323 0.340 0.357 0.385 0.381 0.416 0.415 0.405 0.397 0.359 0.370 0.365 0.350
Avg 0.215 0.261 0.222 0.269 0.227 0.273 0.234 0.270 0.219 0.267 0.265 0.305 0.292 0.313 0.288 0.309 0.256 0.289 0.265 0.307 0.260 0.275
Electricity
96 0.206 0.294 0.170 0.269 0.161 0.261 0.132 0.229 0.181 0.272 0.209 0.307 0.210 0.312 - - - - 0.166 0.263 0.152 0.242
192 0.247 0.338 0.198 0.298 0.190 0.292 0.152 0.250 0.195 0.284 0.219 0.315 0.239 0.337 - - - - 0.191 0.286 0.171 0.259
336 0.310 0.398 0.230 0.331 0.227 0.326 0.173 0.271 0.253 0.338 0.236 0.330 0.284 0.372 - - - - 0.237 0.336 0.192 0.278
720 0.383 0.455 0.288 0.379 0.289 0.372 0.218 0.311 0.304 0.382 0.273 0.328 0.456 0.479 - - - - 0.292 0.384 0.236 0.313
Avg 0.287 0.371 0.222 0.319 0.217 0.313 0.169 0.265 0.233 0.319 0.234 0.320 0.297 0.375 - - - - 0.222 0.317 0.188 0.273
Traffic
96 0.591 0.401 0.522 0.360 0.480 0.344 - - 0.525 0.356 0.572 0.407 0.526 0.368 0.562 0.378 - - 0.514 0.347 - -
192 0.623 0.431 0.538 0.374 0.511 0.369 - - 0.547 0.365 0.575 0.406 0.561 0.385 0.579 0.412 - - 0.543 0.373 - -
336 0.697 0.485 0.574 0.400 0.553 0.394 - - 0.645 0.419 0.588 0.411 0.614 0.412 0.594 0.420 - - 0.575 0.389 - -
720 0.782 0.528 0.669 0.449 0.639 0.429 - - 0.722 0.457 0.618 0.422 0.749 0.464 0.723 0.472 - - 0.622 0.433 - -
Avg 0.673 0.461 0.576 0.396 0.546 0.384 - - 0.610 0.399 0.588 0.412 0.613 0.407 0.615 0.421 - - 0.564 0.386 - -
Solar
96 0.185 0.263 0.167 0.259 0.153 0.255 0.204 0.230 0.209 0.309 0.524 0.557 0.591 0.504 0.373 0.304 0.304 0.345 0.863 0.664 0.682 0.688
192 0.201 0.276 0.180 0.273 0.165 0.269 0.221 0.248 0.220 0.315 0.507 0.550 0.689 0.567 0.363 0.303 0.309 0.342 0.823 0.695 0.694 0.695
336 0.212 0.283 0.190 0.282 0.174 0.278 0.225 0.260 0.228 0.314 0.508 0.553 0.831 0.636 0.391 0.319 0.433 0.450 0.835 0.741 0.719 0.706
720 0.213 0.291 0.198 0.295 0.179 0.287 0.233 0.272 0.218 0.283 0.479 0.534 0.972 0.710 0.444 0.349 0.599 0.576 0.738 0.738 0.759 0.725
Avg 0.203 0.278 0.184 0.277 0.168 0.272 0.221 0.252 0.219 0.305 0.505 0.549 0.771 0.604 0.393 0.319 0.411 0.428 0.815 0.710 0.714 0.704
PEMS08
96 0.249 0.357 0.241 0.353 0.224 0.331 - - 0.608 0.562 1.373 0.995 0.625 0.580 1.538 0.983 - - 1.284 0.888 - -
192 0.299 0.375 0.288 0.368 0.273 0.342 - - 0.774 0.648 1.365 0.979 0.798 0.661 1.719 0.983 - - 1.638 1.039 - -
336 0.330 0.376 0.324 0.367 0.305 0.348 - - 0.892 0.693 1.338 0.960 0.910 0.716 1.768 0.998 - - 1.979 1.160 - -
720 0.418 0.428 0.395 0.406 0.368 0.384 - - 0.975 0.864 1.401 0.980 1.131 0.824 1.802 1.133 - - 2.020 1.175 - -
Avg 0.324 0.384 0.312 0.374 0.293 0.351 - - 0.812 0.692 1.369 0.979 0.866 0.695 1.707 1.024 - - 1.730 1.066 - -
Wind
96 0.948 0.663 0.906 0.650 0.867 0.634 0.931 0.646 0.988 0.678 1.072 0.747 0.946 0.659 1.273 0.738 - - 1.077 0.701 0.992 0.656
192 1.107 0.746 1.062 0.735 1.005 0.714 1.136 0.751 1.275 0.862 1.209 0.804 1.142 0.758 1.439 0.817 - - 1.350 0.825 1.221 0.765
336 1.215 0.801 1.208 0.801 1.097 0.765 1.285 0.820 1.332 0.897 1.318 0.848 1.300 0.830 1.550 0.869 - - 1.473 0.891 1.403 0.844
720 1.307 0.849 1.303 0.847 1.197 0.817 1.390 0.870 1.572 0.906 1.404 0.881 1.417 0.884 1.649 0.914 - - 1.447 0.899 1.581 0.915
Avg 1.144 0.765 1.120 0.758 1.042 0.733 1.186 0.772 1.292 0.836 1.251 0.820 1.201 0.783 1.478 0.834 - - 1.337 0.829 1.299 0.795
NYSE
24 0.323 0.429 0.240 0.327 0.301 0.385 0.370 0.393 0.362 0.403 - - 0.388 0.409 0.480 0.449 - - - - - -
36 0.451 0.520 0.361 0.391 0.437 0.461 0.681 0.566 0.574 0.592 - - 0.778 0.613 0.912 0.655 - - - - - -
48 0.571 0.593 0.487 0.446 0.589 0.538 1.043 0.729 0.688 0.708 - - 1.200 0.792 1.371 0.818 - - - - - -
60 0.762 0.671 0.681 0.557 0.789 0.644 1.425 0.880 0.855 0.802 - - 1.584 1.000 1.754 0.958 - - - - - -
Avg 0.527 0.553 0.442 0.430 0.529 0.507 0.880 0.642 0.620 0.626 - - 0.988 0.704 1.129 0.720 - - - - - -
1st Count 8 8 6 4 15 9 9 9 0 3 3 2 0 0 0 0 1 2 2 3 0 4
2nd Count 3 1 15 2 10 9 2 7 8 6 2 3 0 0 0 0 2 1 5 3 4 4
We provide all the results of the zero-shot probabilistic forecasting in Table 14. As shown in Table 14, we include 9 representative real-world datasets from ProbTS, demonstrating that FLAME achieves state-of-the-art forecasting performance.
Table 14: Full results of zero-shot probabilistic forecasting experiments on datasets from ProbTS. Lower CRPS or NMAE values indicate better predictions. (’-’) denotes datasets included in the model’s pretraining and therefore excluded from testing. (’/’) denotes the excessive time consumption. Red: the best, Blue: the 2nd best.
Models FLAME Small FLAME Base FLAME Large Sundial Chronos MOIRAI Lag-Llama TSDiff CSDI TimeGrad GRU NVP
(Ours) (Ours) (Ours) (2025) (2024) (2024) (2023) (2023) (2022) (2022) (2021)
Metrics CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE
96 0.294 0.368 0.326 0.417 0.315 0.407 0.253 0.308 0.360 0.422 0.464 0.522 0.354 0.402 0.344 0.441 0.236 0.308 0.522 0.645 0.383 0.488
192 0.332 0.420 0.363 0.469 0.337 0.426 0.279 0.337 0.404 0.450 0.467 0.531 0.368 0.415 0.345 0.441 0.291 0.377 0.603 0.748 0.396 0.514
336 0.341 0.427 0.393 0.511 0.351 0.447 0.291 0.350 0.425 0.456 0.524 0.558 0.387 0.436 0.462 0.571 0.322 0.419 0.601 0.759 0.486 0.630
720 0.364 0.454 0.477 0.622 0.383 0.476 0.318 0.380 0.461 0.478 0.514 0.535 0.403 0.466 0.478 0.622 0.448 0.578 0.621 0.793 0.546 0.707
ETTm1
avg 0.333 0.417 0.390 0.505 0.347 0.439 0.285 0.344 0.413 0.451 0.492 0.536 0.378 0.430 0.407 0.519 0.324 0.421 0.587 0.736 0.453 0.585
96 0.120 0.153 0.121 0.153 0.124 0.155 0.128 0.153 0.134 0.158 0.176 0.186 0.163 0.192 0.175 0.224 0.115 0.146 0.427 0.525 0.319 0.413
192 0.146 0.182 0.147 0.184 0.143 0.175 0.150 0.177 0.163 0.183 0.197 0.207 0.181 0.207 0.255 0.316 0.147 0.189 0.424 0.530 0.326 0.427
336 0.163 0.202 0.164 0.205 0.158 0.193 0.167 0.195 0.190 0.204 0.229 0.227 0.206 0.229 0.328 0.397 0.190 0.248 0.469 0.566 0.449 0.580
720 0.185 0.223 0.186 0.231 0.184 0.219 0.189 0.217 0.223 0.230 0.321 0.258 0.227 0.249 0.344 0.416 0.239 0.306 0.470 0.561 0.561 0.749
ETTm2
avg 0.154 0.190 0.155 0.193 0.152 0.186 0.158 0.186 0.178 0.194 0.231 0.220 0.194 0.219 0.276 0.338 0.173 0.222 0.448 0.546 0.414 0.542
96 0.265 0.322 0.261 0.319 0.259 0.310 0.260 0.307 0.293 0.341 0.469 0.488 0.297 0.340 0.395 0.510 0.437 0.557 0.455 0.585 0.379 0.481
192 0.288 0.346 0.282 0.351 0.278 0.342 0.297 0.347 0.348 0.384 0.496 0.492 0.312 0.356 0.467 0.596 0.496 0.625 0.516 0.680 0.425 0.531
336 0.300 0.371 0.297 0.368 0.290 0.362 0.314 0.365 0.384 0.409 0.503 0.489 0.326 0.370 0.450 0.581 0.454 0.574 0.512 0.666 0.458 0.580
720 0.320 0.402 0.312 0.386 0.303 0.367 0.312 0.366 0.413 0.425 0.526 0.508 0.340 0.405 0.516 0.657 0.528 0.657 0.523 0.672 0.502 0.643
ETTh1
avg 0.293 0.360 0.288 0.356 0.283 0.345 0.296 0.346 0.360 0.390 0.498 0.494 0.319 0.368 0.457 0.586 0.479 0.603 0.502 0.651 0.441 0.559
96 0.139 0.177 0.145 0.185 0.144 0.186 0.157 0.186 0.163 0.191 0.212 0.238 0.178 0.204 0.336 0.421 0.164 0.214 0.358 0.448 0.432 0.548
192 0.152 0.195 0.161 0.207 0.159 0.204 0.183 0.213 0.195 0.218 0.229 0.250 0.193 0.217 0.265 0.339 0.226 0.294 0.457 0.575 0.625 0.766
336 0.170 0.216 0.174 0.224 0.172 0.217 0.203 0.232 0.226 0.241 0.253 0.263 0.211 0.230 0.350 0.427 0.274 0.353 0.481 0.606 0.793 0.942
720 0.183 0.231 0.191 0.241 0.182 0.228 0.203 0.233 0.255 0.263 0.279 0.273 0.222 0.238 0.406 0.482 0.302 0.382 0.445 0.550 0.539 0.688
ETTh2
avg 0.161 0.205 0.168 0.214 0.164 0.209 0.186 0.216 0.210 0.228 0.243 0.256 0.201 0.222 0.339 0.417 0.242 0.311 0.435 0.545 0.597 0.736
96 0.071 0.088 0.071 0.085 0.067 0.083 0.078 0.091 0.136 0.155 0.161 0.141 0.085 0.094 0.104 0.113 0.068 0.087 0.130 0.164 0.116 0.145
192 0.089 0.113 0.081 0.100 0.068 0.082 0.083 0.097 0.145 0.165 0.173 0.139 0.094 0.102 0.134 0.144 0.068 0.086 0.127 0.158 0.122 0.147
336 0.094 0.121 0.090 0.110 0.079 0.094 0.090 0.106 0.136 0.151 0.149 0.134 0.098 0.109 0.137 0.138 0.083 0.098 0.130 0.162 0.128 0.160
720 0.103 0.132 0.106 0.125 0.083 0.100 0.096 0.113 0.151 0.161 0.233 0.155 0.107 0.120 0.152 0.141 0.087 0.102 0.113 0.136 0.110 0.135
Weather
avg 0.089 0.114 0.087 0.105 0.074 0.090 0.087 0.102 0.142 0.158 0.179 0.143 0.096 0.106 0.132 0.134 0.077 0.093 0.125 0.155 0.119 0.147
96 0.071 0.090 0.066 0.084 0.069 0.089 0.068 0.083 - - 0.231 0.275 - - 0.344 0.441 0.153 0.203 0.096 0.119 0.094 0.118
192 0.086 0.103 0.076 0.098 0.072 0.088 0.074 0.090 - - 0.236 0.279 - - 0.345 0.441 0.200 0.264 0.100 0.124 0.097 0.121
336 0.096 0.125 0.091 0.117 0.081 0.096 0.083 0.100 - - 0.245 0.289 - - 0.462 0.571 / / 0.102 0.126 0.099 0.123
720 0.117 0.154 0.110 0.141 0.100 0.116 0.098 0.118 - - 0.274 0.318 - - 0.478 0.622 / / 0.108 0.134 0.114 0.144
Electricity
avg 0.093 0.118 0.086 0.110 0.081 0.097 0.081 0.098 - - 0.247 0.290 - - 0.407 0.519 / / 0.102 0.126 0.101 0.127
96 0.240 0.293 0.229 0.277 0.220 0.268 - - 0.250 0.289 - - 0.255 0.297 0.294 0.342 / / 0.202 0.234 0.187 0.231
192 0.257 0.319 0.239 0.291 0.238 0.294 - - 0.249 0.278 - - 0.295 0.343 0.306 0.354 / / 0.208 0.239 0.192 0.236
336 0.271 0.341 0.250 0.311 0.254 0.319 - - 0.269 0.290 - - 0.335 0.384 0.317 0.392 / / 0.213 0.246 0.201 0.248
720 0.304 0.395 0.274 0.346 0.276 0.351 - - 0.310 0.321 - - 0.434 0.517 0.391 0.478 / / 0.220 0.263 0.211 0.264
Traffic
avg 0.268 0.337 0.248 0.306 0.247 0.308 - - 0.269 0.295 - - 0.330 0.385 0.327 0.392 / / 0.211 0.246 0.198 0.245
96 0.023 0.028 0.021 0.027 0.020 0.026 0.024 0.027 0.021 0.025 0.025 0.030 0.042 0.051 0.079 0.090 0.028 0.036 0.068 0.079 0.071 0.091
192 0.032 0.037 0.030 0.036 0.031 0.035 0.034 0.038 0.032 0.036 0.034 0.039 0.047 0.058 0.093 0.106 0.045 0.058 0.087 0.100 0.068 0.087
336 0.044 0.049 0.041 0.047 0.044 0.046 0.046 0.050 0.045 0.048 0.047 0.052 0.061 0.073 0.081 0.106 0.060 0.076 0.074 0.086 0.072 0.091
720 0.074 0.079 0.069 0.077 0.073 0.079 0.076 0.079 0.078 0.080 0.073 0.081 0.078 0.094 0.082 0.142 0.143 0.173 0.099 0.113 0.079 0.103
Exchange
avg 0.043 0.048 0.040 0.047 0.042 0.047 0.045 0.049 0.044 0.047 0.045 0.050 0.057 0.069 0.084 0.111 0.069 0.086 0.082 0.095 0.073 0.093
24 0.135 0.166 0.111 0.143 0.136 0.168 0.108 0.123 0.120 0.139 0.150 0.196 0.135 0.173 0.228 0.242 0.250 0.263 0.275 0.296 0.257 0.283
36 0.164 0.200 0.119 0.153 0.145 0.179 0.152 0.169 0.179 0.205 0.171 0.222 0.163 0.227 0.235 0.246 0.285 0.298 0.272 0.298 0.281 0.307
48 0.176 0.216 0.139 0.178 0.164 0.200 0.164 0.184 0.186 0.215 0.151 0.184 0.171 0.233 0.265 0.275 0.285 0.301 0.295 0.320 0.288 0.314
60 0.171 0.211 0.149 0.189 0.169 0.204 0.168 0.190 0.196 0.228 0.163 0.188 0.156 0.211 0.263 0.272 0.283 0.299 0.295 0.325 0.307 0.333
ILI
avg 0.162 0.198 0.130 0.166 0.154 0.188 0.148 0.166 0.170 0.197 0.159 0.197 0.156 0.211 0.248 0.259 0.276 0.290 0.284 0.310 0.283 0.309
1st Count 3 3 7 3 15 14 5 8 0 1 0 1 0 0 0 0 2 2 0 2 4 2
2nd Count 7 5 5 5 7 6 5 8 0 2 1 1 1 0 0 0 6 5 4 2 0 2
Considering short-term forecasting, we also conduct experiments on 8,068 univariate datasets in TFB (Qiu et al., 2024), we report the average MASE and msMAPE results in Figures 6 and 7. FLAME Large outperforms all baselines including end-to-end supervised models and foundation models.
Figure 6: The average MASE results of 8068 datasets in TFB.
Figure 7: The average msMAPE results of 8068 datasets in TFB.
C.2 Full-shot and few-shot results
We further compare FLAME Small with end-to-end supervised models. We provide all the results of deterministic forecasting in Table 15, comparing FLAME Small of different training budgets with most advanced end-to-end determinsitic forecasting models.
We also provide all the results of probabilistic forecasting in Table 16, comparing FLAME Small of different training budgets with most advanced end-to-end probabilistic forecasting models.
Table 15: Deterministic forecasting results. Full-shot, 10% few-shot and zero-shot results of FLAME Small v.s most advanced end-to-end supervised models on commonly-used datasets in TSFM-Bench. Lower MSE or MAE values indicate better predictions. Red: the best, Blue: the second best.
Models FLAME Small FLAME Small FLAME Small TimeKAN AMD TimePro TimeXer Fredformer iTransformer PatchTST TimesNet
(Zero-shot) (10% few-shot)) (Full-shot) (2025) (2025) (2025) (2024) (2024) (2024) (2023) (2023)
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTm1
96 0.305 0.353 0.278 0.328 0.264 0.318 0.327 0.365 0.327 0.361 0.326 0.364 0.309 0.352 0.326 0.361 0.334 0.368 0.329 0.367 0.338 0.375
192 0.328 0.370 0.320 0.359 0.311 0.352 0.363 0.387 0.366 0.383 0.367 0.383 0.355 0.378 0.363 0.384 0.377 0.391 0.367 0.385 0.374 0.387
336 0.349 0.384 0.338 0.371 0.332 0.368 0.389 0.407 0.398 0.404 0.402 0.409 0.387 0.399 0.395 0.406 0.426 0.420 0.399 0.410 0.410 0.411
720 0.392 0.416 0.383 0.404 0.372 0.399 0.457 0.445 0.464 0.437 0.469 0.446 0.448 0.435 0.456 0.441 0.491 0.459 0.454 0.439 0.478 0.450
Avg 0.344 0.381 0.330 0.366 0.320 0.359 0.384 0.401 0.389 0.396 0.391 0.401 0.375 0.391 0.385 0.398 0.407 0.410 0.387 0.400 0.400 0.406
ETTm2
96 0.165 0.255 0.155 0.250 0.150 0.243 0.178 0.262 0.176 0.259 0.178 0.260 0.171 0.255 0.177 0.258 0.180 0.264 0.175 0.259 0.187 0.267
192 0.217 0.297 0.215 0.294 0.214 0.292 0.244 0.308 0.242 0.302 0.242 0.303 0.238 0.300 0.243 0.301 0.250 0.309 0.241 0.302 0.249 0.309
336 0.262 0.330 0.258 0.318 0.253 0.312 0.305 0.346 0.298 0.337 0.303 0.342 0.301 0.340 0.302 0.340 0.311 0.348 0.305 0.343 0.321 0.351
720 0.333 0.379 0.329 0.368 0.328 0.364 0.402 0.404 0.396 0.394 0.400 0.399 0.401 0.397 0.404 0.398 0.412 0.407 0.402 0.400 0.408 0.403
Avg 0.244 0.315 0.239 0.308 0.236 0.303 0.282 0.330 0.278 0.323 0.281 0.326 0.278 0.323 0.281 0.324 0.288 0.332 0.281 0.326 0.291 0.333
ETTh1
96 0.362 0.395 0.357 0.386 0.342 0.375 0.374 0.391 0.375 0.392 0.375 0.398 0.377 0.397 0.378 0.395 0.386 0.405 0.414 0.419 0.384 0.402
192 0.406 0.427 0.392 0.413 0.384 0.404 0.421 0.421 0.430 0.422 0.427 0.429 0.425 0.425 0.435 0.424 0.441 0.436 0.460 0.445 0.436 0.429
336 0.436 0.449 0.420 0.435 0.407 0.428 0.464 0.440 0.471 0.443 0.472 0.450 0.457 0.441 0.485 0.447 0.487 0.458 0.501 0.466 0.491 0.469
720 0.471 0.482 0.436 0.450 0.418 0.436 0.466 0.462 0.478 0.464 0.476 0.474 0.464 0.463 0.496 0.472 0.503 0.491 0.500 0.488 0.521 0.500
Avg 0.419 0.438 0.401 0.421 0.388 0.411 0.431 0.429 0.438 0.430 0.438 0.438 0.431 0.432 0.448 0.435 0.454 0.448 0.469 0.455 0.458 0.450
ETTh2
96 0.278 0.339 0.271 0.327 0.264 0.320 0.293 0.343 0.287 0.338 0.293 0.345 0.289 0.340 0.291 0.342 0.297 0.349 0.302 0.348 0.340 0.374
192 0.336 0.381 0.335 0.376 0.332 0.364 0.375 0.396 0.367 0.388 0.367 0.394 0.370 0.391 0.372 0.390 0.380 0.400 0.388 0.400 0.402 0.414
336 0.373 0.410 0.364 0.402 0.360 0.399 0.429 0.441 0.410 0.424 0.419 0.431 0.422 0.434 0.419 0.431 0.428 0.432 0.426 0.433 0.452 0.452
720 0.411 0.441 0.391 0.428 0.382 0.423 0.466 0.468 0.421 0.440 0.427 0.445 0.429 0.445 0.431 0.450 0.427 0.445 0.431 0.446 0.462 0.468
Avg 0.350 0.393 0.340 0.383 0.335 0.377 0.391 0.412 0.371 0.397 0.377 0.404 0.378 0.403 0.378 0.403 0.383 0.407 0.387 0.407 0.414 0.427
Weather
96 0.148 0.204 0.144 0.197 0.141 0.184 0.164 0.210 0.174 0.221 0.166 0.207 0.168 0.209 0.163 0.207 0.174 0.214 0.177 0.218 0.172 0.220
192 0.189 0.243 0.184 0.241 0.182 0.240 0.209 0.250 0.219 0.259 0.216 0.254 0.220 0.254 0.224 0.258 0.221 0.254 0.255 0.259 0.219 0.261
336 0.233 0.278 0.230 0.276 0.224 0.272 0.264 0.290 0.273 0.296 0.273 0.296 0.276 0.296 0.278 0.298 0.278 0.296 0.278 0.297 0.280 0.306
720 0.289 0.320 0.287 0.318 0.280 0.306 0.343 0.342 0.349 0.345 0.351 0.346 0.353 0.347 0.357 0.350 0.358 0.349 0.354 0.348 0.365 0.359
Avg 0.215 0.261 0.211 0.258 0.207 0.251 0.245 0.273 0.254 0.280 0.252 0.276 0.254 0.277 0.256 0.278 0.258 0.278 0.266 0.281 0.259 0.287
Electricity
96 0.206 0.294 0.142 0.235 0.124 0.218 0.174 0.266 0.147 0.251 0.139 0.234 0.151 0.247 0.148 0.242 0.148 0.240 0.195 0.285 0.168 0.272
192 0.247 0.338 0.168 0.255 0.149 0.236 0.182 0.273 0.176 0.262 0.156 0.249 0.165 0.261 0.165 0.257 0.162 0.253 0.199 0.289 0.184 0.289
336 0.310 0.398 0.188 0.271 0.164 0.252 0.197 0.286 0.193 0.281 0.172 0.267 0.183 0.280 0.180 0.274 0.178 0.269 0.215 0.305 0.198 0.300
720 0.383 0.455 0.192 0.277 0.205 0.288 0.236 0.320 0.232 0.329 0.209 0.299 0.220 0.309 0.218 0.305 0.225 0.317 0.256 0.337 0.220 0.320
Avg 0.287 0.371 0.173 0.260 0.161 0.249 0.197 0.286 0.187 0.281 0.169 0.262 0.180 0.274 0.178 0.270 0.178 0.270 0.216 0.304 0.193 0.295
Traffic
96 0.591 0.401 0.375 0.262 0.358 0.246 0.423 0.286 0.443 0.298 0.426 0.292 0.416 0.280 0.403 0.274 0.395 0.268 0.544 0.359 0.593 0.321
192 0.623 0.431 0.388 0.265 0.379 0.248 0.442 0.295 0.496 0.323 0.439 0.298 0.435 0.288 0.429 0.289 0.417 0.276 0.540 0.354 0.617 0.336
336 0.697 0.485 0.397 0.285 0.386 0.262 0.473 0.335 0.520 0.330 0.449 0.307 0.451 0.296 0.441 0.295 0.433 0.283 0.551 0.358 0.629 0.336
720 0.782 0.528 0.451 0.332 0.427 0.282 0.481 0.357 0.540 0.344 0.475 0.309 0.484 0.314 0.463 0.300 0.467 0.302 0.586 0.375 0.640 0.350
Avg 0.673 0.461 0.403 0.286 0.388 0.260 0.455 0.318 0.500 0.324 0.447 0.302 0.447 0.295 0.434 0.289 0.428 0.282 0.555 0.362 0.620 0.336
1st Count 0 0 1 1 34 34 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
Table 16: Probabilistic forecasting results. Full-shot, 10% few-shot and zero-shot results of FLAME Small v.s most advanced end-to-end supervised models on commonly-used datasets in ProbTS. Lower CRPS or NMAE values indicate better predictions. Red: the best, Blue: the second best.
Models FLAME Small FLAME Small FLAME Small NSDIFF $K^{2}$VAE TSDiff CSDI TimeGrad GRU NVP
(Zero-shot) (10% few-shot) (Full-shot) (2025) (2025) (2023) (2022) (2022) (2021)
Metric CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE
ETTm1
96 0.294 0.368 0.233 0.287 0.211 0.275 0.295 0.364 0.235 0.294 0.344 0.441 0.236 0.308 0.522 0.645 0.383 0.488
192 0.332 0.420 0.259 0.317 0.232 0.298 0.285 0.350 0.261 0.330 0.345 0.441 0.291 0.377 0.603 0.748 0.396 0.514
336 0.341 0.427 0.292 0.351 0.279 0.355 0.289 0.357 0.304 0.386 0.462 0.571 0.322 0.419 0.601 0.759 0.486 0.630
720 0.364 0.454 0.304 0.392 0.282 0.358 0.411 0.500 0.302 0.387 0.478 0.622 0.448 0.578 0.621 0.793 0.546 0.707
Avg 0.333 0.417 0.272 0.337 0.251 0.322 0.320 0.393 0.276 0.349 0.407 0.519 0.324 0.421 0.587 0.736 0.453 0.585
ETTm2
96 0.120 0.153 0.117 0.144 0.114 0.141 0.145 0.176 0.128 0.148 0.175 0.224 0.115 0.146 0.427 0.525 0.319 0.413
192 0.146 0.182 0.145 0.178 0.138 0.166 0.162 0.196 0.161 0.163 0.255 0.316 0.147 0.189 0.424 0.530 0.326 0.427
336 0.163 0.202 0.160 0.199 0.151 0.182 0.172 0.209 0.180 0.183 0.328 0.397 0.190 0.248 0.469 0.566 0.449 0.580
720 0.185 0.223 0.181 0.231 0.170 0.211 0.364 0.428 0.224 0.224 0.344 0.416 0.239 0.306 0.470 0.561 0.561 0.749
Avg 0.154 0.190 0.151 0.188 0.143 0.175 0.211 0.252 0.173 0.180 0.276 0.338 0.173 0.222 0.448 0.546 0.414 0.542
ETTh1
96 0.265 0.322 0.254 0.311 0.247 0.301 0.301 0.380 0.280 0.358 0.395 0.510 0.437 0.557 0.455 0.585 0.379 0.481
192 0.288 0.346 0.272 0.336 0.268 0.322 0.309 0.386 0.304 0.392 0.467 0.596 0.496 0.625 0.516 0.680 0.425 0.531
336 0.300 0.371 0.297 0.369 0.288 0.363 0.357 0.448 0.326 0.413 0.450 0.581 0.454 0.574 0.512 0.666 0.458 0.580
720 0.320 0.402 0.308 0.395 0.300 0.392 0.384 0.513 0.331 0.428 0.516 0.657 0.528 0.657 0.523 0.672 0.502 0.643
Avg 0.293 0.360 0.283 0.353 0.276 0.345 0.337 0.432 0.310 0.398 0.457 0.586 0.479 0.603 0.502 0.651 0.441 0.559
ETTh2
96 0.139 0.177 0.133 0.175 0.122 0.162 0.175 0.220 0.182 0.206 0.336 0.421 0.164 0.214 0.358 0.448 0.432 0.548
192 0.152 0.195 0.148 0.193 0.137 0.188 0.200 0.245 0.185 0.220 0.265 0.339 0.226 0.294 0.457 0.575 0.625 0.766
336 0.170 0.216 0.167 0.209 0.155 0.194 0.207 0.257 0.245 0.277 0.350 0.427 0.274 0.353 0.481 0.606 0.793 0.942
720 0.183 0.231 0.177 0.228 0.171 0.222 0.209 0.260 0.237 0.272 0.406 0.482 0.302 0.382 0.445 0.550 0.539 0.688
Avg 0.161 0.205 0.156 0.201 0.146 0.192 0.198 0.246 0.212 0.244 0.339 0.417 0.242 0.311 0.435 0.545 0.597 0.736
Weather
96 0.071 0.088 0.069 0.085 0.065 0.083 0.123 0.129 0.083 0.087 0.104 0.113 0.068 0.087 0.130 0.164 0.116 0.145
192 0.089 0.113 0.078 0.101 0.073 0.088 0.132 0.132 0.085 0.090 0.134 0.144 0.068 0.086 0.127 0.158 0.122 0.147
336 0.094 0.121 0.082 0.114 0.075 0.089 0.130 0.128 0.087 0.091 0.137 0.138 0.083 0.098 0.130 0.162 0.128 0.160
720 0.103 0.132 0.098 0.119 0.091 0.096 0.142 0.147 0.090 0.094 0.152 0.141 0.087 0.102 0.113 0.136 0.110 0.135
Avg 0.089 0.114 0.082 0.105 0.076 0.089 0.132 0.134 0.086 0.091 0.132 0.134 0.077 0.093 0.125 0.155 0.119 0.147
Electricity
96 0.071 0.090 0.069 0.088 0.069 0.088 0.160 0.202 0.089 0.093 0.344 0.441 0.153 0.203 0.096 0.119 0.094 0.118
192 0.086 0.103 0.083 0.097 0.084 0.095 0.248 0.358 0.097 0.102 0.345 0.441 0.200 0.264 0.100 0.124 0.097 0.121
336 0.096 0.125 0.097 0.116 0.094 0.114 0.252 0.362 0.101 0.105 0.462 0.571 / / 0.102 0.126 0.099 0.123
720 0.117 0.154 0.108 0.123 0.104 0.117 0.259 0.368 0.105 0.110 0.478 0.622 / / 0.108 0.134 0.114 0.144
Avg 0.093 0.118 0.089 0.106 0.088 0.104 0.230 0.322 0.098 0.103 0.407 0.519 / / 0.102 0.126 0.101 0.127
Traffic
96 0.240 0.293 0.227 0.262 0.203 0.217 0.393 0.500 0.216 0.222 0.294 0.342 / / 0.202 0.234 0.187 0.231
192 0.257 0.319 0.243 0.289 0.212 0.223 0.448 0.585 0.220 0.227 0.306 0.354 / / 0.208 0.239 0.192 0.236
336 0.271 0.341 0.266 0.317 0.220 0.228 0.455 0.592 0.228 0.235 0.317 0.392 / / 0.213 0.246 0.201 0.248
720 0.304 0.395 0.288 0.322 0.231 0.237 0.452 0.588 0.235 0.242 0.391 0.478 / / 0.220 0.263 0.211 0.264
Avg 0.268 0.337 0.256 0.298 0.217 0.226 0.437 0.566 0.225 0.232 0.327 0.392 / / 0.211 0.246 0.198 0.245
Exchange
96 0.023 0.028 0.024 0.031 0.022 0.027 0.023 0.029 0.027 0.028 0.079 0.090 0.028 0.036 0.068 0.079 0.071 0.091
192 0.032 0.037 0.028 0.036 0.025 0.031 0.029 0.037 0.035 0.036 0.093 0.106 0.045 0.058 0.087 0.100 0.068 0.087
336 0.044 0.049 0.035 0.037 0.031 0.033 0.043 0.053 0.027 0.028 0.081 0.106 0.060 0.076 0.074 0.086 0.072 0.091
720 0.074 0.079 0.067 0.074 0.053 0.070 0.057 0.071 0.081 0.083 0.082 0.142 0.143 0.173 0.099 0.113 0.079 0.103
Avg 0.043 0.048 0.039 0.045 0.033 0.040 0.038 0.047 0.043 0.044 0.084 0.111 0.069 0.086 0.082 0.095 0.073 0.093
ILI
24 0.135 0.166 0.112 0.133 0.088 0.103 0.092 0.108 0.110 0.113 0.228 0.242 0.250 0.263 0.275 0.296 0.257 0.283
36 0.164 0.200 0.148 0.185 0.137 0.168 0.151 0.178 0.140 0.144 0.235 0.246 0.285 0.298 0.272 0.298 0.281 0.307
48 0.176 0.216 0.169 0.205 0.142 0.171 0.151 0.178 0.153 0.157 0.265 0.275 0.285 0.301 0.295 0.320 0.288 0.314
60 0.171 0.211 0.165 0.198 0.152 0.184 0.168 0.198 0.164 0.167 0.263 0.272 0.283 0.299 0.295 0.325 0.307 0.333
Avg 0.162 0.198 0.149 0.180 0.130 0.157 0.140 0.165 0.142 0.145 0.248 0.259 0.276 0.290 0.284 0.310 0.283 0.309
1st Count 0 0 1 1 28 26 0 0 1 8 0 0 2 1 0 0 4 0
Appendix D Showcases
We visualized the forecasting results of FLAME Large on TSFM-Bench and TFB in Figures 8 and 9.
(a) Electricity
(b) ETTh1
(c) ETTh2
(d) ETTm1
(e) ETTm2
(f) PEMS08
(g) Solar
(h) Traffic
(i) Weather
(j) Wind
Figure 8: Visualization of TSFM-Bench
(a) economics_36
(b) human_58
(c) m3_monthly_dataset_223
(d) kdd_cup_2018_dataset_without_missing_values_60
(e) m3_other_dataset_7
(f) m4_quarterly_dataset_18009
(g) m4_monthly_dataset_2019
(h) m4_monthly_dataset_84
(i) vehicle_trips_dataset_without_missing_values_82
(j) cif_2016_dataset_65
Figure 9: Visualization of Univariate Datasets in TFB
Appendix E Related Works
Time Series Analysis holds a position of paramount importance across a diverse range of fields, including the economy (Qiu et al., 2025b, a; Liu et al., 2025a), transportation (Wu et al., 2025d, 2024), health (Lu et al., 2023, 2024a, 2024b), weather (Li et al., 2025a; Yang et al., 2024; Zhou et al., 2025), and energy (Sun et al., 2025a). It encompasses a multitude of critical tasks, such as forecasting (Li et al., 2025b; Dai et al., 2024c), anomaly detection (Wang et al., 2025a; Qiu et al., 2025c; Wu et al., 2025c; Zhang et al., 2025), classification (Liu et al., 2024b), imputation (Wu et al., 2023), and others (Qiu et al., 2025f; Wu et al., 2025b). Among these tasks, Time Series Forecasting is the one most extensively applied in practical, real-world scenarios.
Time series forecasting (TSF) involves the prediction of future observations on the basis of historical data. Research results have demonstrated that features learned through specific methods may present superior performance in comparison to human-designed features (Dai et al., 2024b; Hu et al., 2025a; Dai et al., 2024c; Liu et al., 2025c; Sun et al., 2025b, a; Niu et al., 2025; Qiu et al., 2025d). By leveraging the representation learning capabilities of deep neural networks (DNNs), numerous deep-learning based approaches have come into existence. For example, TimesNet (Wu et al., 2023) and SegRNN (Lin et al., 2023) model time series as vector sequences, utilizing CNNs or RNNs to capture temporal dependencies. Furthermore, Transformer architectures, including Informer (Zhou et al., 2021), Dsformer (Yu et al., 2023), TimeFilter (Hu et al., 2025b), TimeBridge (Liu et al., 2025b), PDF (Dai et al., 2024a), Triformer (Cirstea et al., 2022), PatchTST (Nie et al., 2023), ROSE (Wang et al., 2025c), LightGTS (Wang et al., 2025b), and MagicScaler (Pan et al., 2023), are able to more precisely capture the complex relationships between time points, thus significantly improving forecasting performance. MLP-based methods, such as DUET (Qiu et al., 2025e), SRSNet (Wu et al., 2025a), APN (Liu et al., 2026), AMD (Hu et al., 2025a), NLinear (Zeng et al., 2023), and DLinear (Zeng et al., 2023), adopt relatively simpler architectures with fewer parameters yet still succeed in achieving highly competitive forecasting accuracy.